### v1

# Measurement-device agnostic quantum tomography

Abstract
Characterization of quantum states and devices is paramount to quantum science and technology. The characterization consists of individual measurements, which are required to be precisely known. A mismatch between actual and assumed constituent measurements limits the accuracy of this characterization. Here, we show that such a mismatch introduces reconstruction artifacts in quantum state tomography. We use these artifacts to detect and quantify the mismatch and gain information about the actual measurement operators. It consequently allows the mitigation of systematic errors in quantum measurement and state preparation.
Introduction - The measurements play a central role in quantum science and technology [1]. However, the implemented measurements often differ from the ideal ones, leading to systematic errors. One can compensate for these systematic errors with the knowledge of the actual measurements. We investigate this problem from the point of view of quantum tomography, one of the most valuable tools in quantum science [2, 3, 4, 5, 6]. It allows the complete characterization of prepared states and implemented quantum circuits, assessing and certifying their quality. Quantum state tomography consists of a suitable set of known measurements on the investigated objects. One can reconstruct the state from the collection of measurement outcomes, the tomogram. Imperfect realization of these measurements causes a mismatch between the assumed and the realized measurement operators. Such a mismatch leads to the wrong interpretation of the acquired tomogram and manifests as a reconstruction artifact.
The true measurement operators can be fully characterized with measurement tomography [7, 8, 9, 10], which relies on perfect knowledge of input probe quantum states. Such a requirement is difficult to meet in the experiments. A self-consistent tomography framework has been introduced to loosen this condition [11]. The framework suggests the existence of a self-calibrating state class, which could reveal, when measured, the initially unknown information about the states and the measurement device. The cases of photon-number-resolving measurement and quantum homodyne tomography have been discussed. Another approach demonstrated for polarization states of light infers a single unknown parameter of the measurement device using maximum likelihood estimation [12]. Finally, assuming the sparsity of the parameter vector and the low rank of the measured density matrix, the dependence on the measurement device can be relaxed [13].
In this Letter, we show that an ensemble of single-qubit states of equal purity distributed quasi-uniformly on the Bloch sphere forms a class of self-calibrating states. This class is convenient in experiments where states are naturally prepared with quasi-equal purity, e.g., in photonics and trapped ions. Advantageously, the precise control, characterization, or even sparsity of such self-calibrating probe states is not required. We show a novel approach to simultaneously calibrating quantum measurements and state preparation. To this end we exploit artifacts in quantum state reconstruction performed on the self-calibrating probe states. In the multi-qubit setting, we can use the method in a fully scalable way to mitigate systematic errors on each local qubit measurement, effectively mitigating errors in the whole system. The method also applies to quantum device tomography thanks to the state-process duality [14, 15, 1], which allows the quantum processes to be treated as states. Our method is not limited to the determination of an initially unknown parameter shared among all measurement operators in the tomographic scheme; it allows the determination of parameters related to each measurement operator individually. The presented scheme provides the characterization of quantum states independently of measurement device imperfections. General concepts are introduced using numerical simulations, and we support our findings with a photonic experiment.
Figure 1: Purity modulation due to the measurement operator mismatch. States are represented as points on the Bloch sphere plotted using Hammer projection, and the purity of their reconstruction is color-coded. Although all probe states were prepared with the same purity, the reconstruction using quantum tomography with mismatched measurement operators yields purity modulation.
Measurement operator mismatch artifact - In quantum state tomography, many copies of an initially unknown state, described by density operator $\hat{\rho}$, are subject to a series of measurements described by measurement operators $\hat{\pi}_{j}$. These measurement operators should be tomographically complete, and their choice influences the performance of the tomography [16, 17, 18, 19]. The resulting collection of measurement outcomes forms the tomogram. The tomogram and the measurement operators are then input into an algorithm, which outputs the density operator. If the actual measurement operators $\hat{\pi}^{\prime}_{j}$ differ from those input into the reconstruction method, artifacts appear. We use the maximum-likelihood method [20], but in Section S4 of the Supplementary Material, we demonstrate that the effect is not limited to the particular reconstruction method. The disagreement between actual and assumed measurement operators causes inconsistency in the tomogram, which results in state-dependent purity loss.
Let us demonstrate the behavior in the simplest scenario of pure probe single-qubit states and projective measurements onto the eigenstates of Pauli operators, called Pauli tomography. The measurement operator is realized by two rotation angles, corresponding to rotation about the $y$- and $z$-axis and subsequent projection onto computational basis state $|0\rangle\langle 0|$. Throughout the article, we use the computational basis $\{|0\rangle,|1\rangle\}$. The following operators describe the constituent measurements
$\hat{\pi}_{j}=\hat{R}_{z}^{\dagger}(\phi_{j})\hat{R}_{y}^{\dagger}(\theta_{j})|0\rangle\langle 0|\hat{R}_{y}(\theta_{j})\hat{R}_{z}(\phi_{j}),$ (1)
where $\hat{R}_{j}(\alpha)=\exp(i\hat{\sigma}_{j}\alpha/2)$ is the rotation generated by the Pauli operator. Then, the corresponding detection probability is given by the Born rule,
$p_{j}=\mathrm{Tr}[\hat{\rho}\hat{\pi}_{j}].$ (2)
The angles $\theta_{j}$ and $\phi_{j}$ for Pauli tomography are provided in Table S2 of the Supplemental Material. Assume that due to experimental imperfections, the actual angles
$\theta_{j}^{\prime}=\theta_{j}+\delta_{j},$ (3)
$\phi_{j}^{\prime}=\phi_{j}+\epsilon_{j},$ (4)
i.e., we introduce systematic additive errors for each measurement operator. The error vector $\vec{\delta}=\{\delta_{j},\epsilon_{j}\}$ parameterizes the measurement operators $\hat{\pi}^{\prime}_{j}=\hat{\pi}_{j}(\vec{\delta})$. We numerically simulate tomograms using these true measurement operators for many pure probe states covering the Bloch sphere quasi-uniformly. Then, we reconstruct them assuming the original unperturbed measurement operators $\{\hat{\pi}_{j}\}$. This mismatch causes the state-dependent purity modulation, shown in Figure 1. We quantify the purity modulation in the ensemble of reconstructed states as
$\Delta P=\max P-\min P,$ (5)
where $P=\mathrm{Tr}\left(\hat{\rho}^{2}\right)$ is the purity of the reconstructed state. The modulation $\Delta P$ scales with the magnitude of parameters $\vec{\delta}$ and thus quantifies the severity of the measurement operator mismatch. The error vector $\vec{\delta}$ determines at which state the artifact manifests maximally. We will leverage these features to discover the vector $\delta$ and, consequently, the true measurement operators.
Artifact mitigation - We minimize the artifact severity, i.e., purity modulation (5), by varying the measurement operators used to reconstruct the probe states. After recording all tomograms of the probe states we used, we optimize our assumption of $\vec{\delta}$ which specifies the measurement operators used in the reconstruction algorithm. Once the assumed parameters match the actual parameters, the artifact is minimized. Since the perturbed measurement operators no longer add the identity operator, we must properly renormalize them in the reconstruction [20].
Due to the state dependency, the Bloch sphere must be sampled densely enough to detect the mismatch. The probe states should cover the Bloch sphere quasi-uniformly and possess uniform purity; otherwise, their full control is not required. This feature is advantageous in practice because state preparation is never perfect in real-world implementation. It distinguishes our approach from the schemes that rely on perfect knowledge of the probe states, like quantum measurement tomography [8].
Generally, the optimization landscape has multiple local minima; therefore, global optimization must be performed. Here, we use the NOMAD implementation of the MADS algorithm [21]. The resulting optimum $\vec{\delta}$ is not unique because any set of measurement operators $\{\hat{U}\hat{\pi}_{j}\hat{U}^{\dagger}\}$, where $\hat{U}$ is a fixed unitary operation, also minimizes the artifact. Each of these satisfactory sets corresponds to different parameters $\vec{\delta}$. These unitary-equivalent solutions could be interpreted as unitary transformations of probe states. Albeit this ambiguity, such a calibration is already applicable in schemes invariant to local unitaries such as random measurement [22] or entanglement certification [23, 24, 25], where only relative orientations of $\hat{\pi}_{j}$ operators are important. Moreover, the unitary operation can be easily determined by additional measurement of two known and non-orthogonal probe states.
Since the artifact is state-dependent, the probe sampling strategy is important for minimization. The sampling with too few probe states could lead to false error parameters $\vec{\delta}_{f}$. We used 30 quasi-uniformly distributed probe states. Such a choice provided a reliable determination of error parameters. We discuss the effects of the number of probe states in detail in the Supplementary material.
To test the method numerically, we first randomly generate ground-truth error parameters $\vec{\delta}$. Each element is normally distributed with zero mean value and standard deviation of 10 deg. Then, we produce tomograms corresponding to pure probe states $\rho_{k}=|\psi_{k}\rangle\langle\psi_{k}|$. For simplicity, we exclude statistical noise from our simulations. Parameters $\vec{\delta}^{\prime}$ assumed in the reconstruction are optimized to minimize the purity modulation $\Delta P$ in the set of reconstructed probe states $\{\hat{\rho}_{k}\}_{k=1,\dots,30}$, concluding the calibration step.
To verify the correctness of the optimization result, we use the ground truth $\vec{\delta}$ again to generate tomograms of pure test states $\hat{\xi}=|\xi_{k}\rangle\langle\xi_{k}|$, which quasi-uniformly covers the Bloch sphere in 108 points. We reconstruct these states using the optimized parameters $\vec{\delta}$ obtained in the calibration step and calculate the purity of the reconstructions. As a further test, we assume knowledge of two non-orthogonal probe states and use their reconstructions to determine corrective unitary operation $\hat{W}$. We apply $\hat{W}$ to all reconstructed test states $\hat{\xi}_{k}$ to eliminate the unitary-related ambiguity. Then, we check their fidelity with the expected ideal test state $F_{k}=\langle\xi_{k}|\hat{\rho}_{k}|\xi_{k}\rangle$. We characterize the test states ensemble with the purity modulation and their lowest fidelity $\min\limits_{k}F_{k}$ to provide the worst-case estimate.
We repeated the whole numerical simulation 100 times, randomly choosing new ground truth for the error parameters $\delta$ in each run. The details about optimizer settings are provided in Section S1 of the Supplementary Material. When we compare the optimized results to the reference, i.e. the states reconstructed with the assumption of null error vector, the purity modulation decreases from $\Delta P=(8\begin{subarray}{c}+3\\ -3\end{subarray}))\cdot 10^{-2}$ to $\Delta P=(1.0\begin{subarray}{c}+0.8\\ -1.0\end{subarray})\cdot 10^{-3}$. The uncertainty interval spans from 0.158 to 0.842-quantile in correspondence to $\pm$ one standard deviation and is used due to the skewed distribution of the results. Mainly, the infidelity decreased from $(3\begin{subarray}{c}+1\\ -2\end{subarray})\cdot 10^{-2}$ to $(3\begin{subarray}{c}+5\\ -3\end{subarray})\cdot 10^{-4}$, which represents two orders of magnitude improvement in accuracy. This improvement is crucial for reaching error levels low enough for quantum error correction schemes. The lower $\Delta P$ of the probe state ensemble was achieved the better results in the test ensemble we achieved. These results illustrate the feasibility of such optimization and its ability to reliably determine the true measurement operators. It is important to stress that the pure probe states and projective (pure) measurements were chosen here for the simplicity of introducing the measurement-device agnostic framework. However, the method is fully applicable in real cases of nonunity purity (as in the following experimental demonstration) and generic positive-operator-valued measurements.
In the multi-qubit scenario, the tomography is usually done using local measurements. In that case, we use the presented method for each local measurement device. The artifact severity scales worst as $P_{\min,n}=\prod\limits_{i=1}^{n}P_{\min,i}$ for $n$-qubit state tomography, where $P_{\min,i}$ is the corresponding minimal purity observed at the single qubit. If the input state |$\ket{\psi_{i}}$ at the $i$-th qubit suffers from the artifact the worst, then product state $\bigotimes\limits_{i}\ket{\psi_{i}}$ is the $n$-qubit state with the greatest manifestation of the artifact. In such a case, the single-qubit improvements multiply and lead to significantly better overall results in multi-qubit scenarios.
State preparation - The state preparation typically suffers from imperfections too. Here, we use the symmetry in the Born rule to formally exchange state preparation and measurement to utilize the previously introduced method to correct state preparation. Usually, the qubit is initialized in state $|0\rangle$, then turned into state $|\psi_{j}\rangle=\hat{U}_{j}|0\rangle$ by unitary evolution $\hat{U}_{j}$. The tomographic projection is also typically realized by some unitary evolution $\hat{V}_{k}$ and subsequent projection, i.e., $|\pi_{k}\rangle=\hat{V}_{k}^{\dagger}|0\rangle$. The state tomogram of the probe state $|\psi_{j}\rangle$ consists of measured probabilities $|\langle\hat{\pi}_{k}|\psi_{j}\rangle|^{2}=|\langle 0|\hat{V}_{k}\hat{U}_{j}|0\rangle|^{2}$. The discussed procedure obtains information related to projections, i.e. about $V_{k}$.
Physical reversal of the process consists of preparing states $\hat{V}_{k}|0\rangle$ and projecting them onto states $\hat{U}_{j}^{\dagger}|0\rangle$, obtaining tomogram elements $p_{kj}\propto|\langle 0|\hat{U}_{j}\hat{V}_{k}|0\rangle|^{2}$. We can analyze the tomograms to improve our knowledge of the measurement operator, i.e., $\hat{U}_{j}^{\dagger}|0\rangle\langle 0|\hat{U}_{j}$, related to state preparation in the non-reversed situation. The physical reversal can be realized in the experiment by performing projections onto the original probe states in the tomographic device and preparing states corresponding to the original tomographical projectors instead of the original probes. Either way, we gain information about $\hat{U}_{j}$, increasing the accuracy of state preparation.
Special case: multiplicative error - An important subclass of the problem is the case of multiplicative measurement operator error
$\theta_{j}^{\prime}=(1+\delta)\theta_{j},$ (6)
$\phi_{j}^{\prime}=(1+\epsilon)\phi_{j}.$ (7)
This case is relevant when we sequentially switch the unitary operations that precede a fixed projector to perform the tomographic measurements. These multiplicative errors could be interpreted as Bloch sphere under- or over-rotation. This approach to quantum state tomography has been reported in various platforms, including optics [26, 27], color-centers in solids [28], trapped ions [29], neutral atoms [30], or superconducting qubits [31], making this problem subclass highly relevant.
Because parameters $(\delta,\epsilon)$ are shared among all projection operators, the optimum search simplifies from 12 parameters to just two. In Section S2 of the Supplementary material, we show that with just eight probe states, the optimization landscape possesses a single minimum, and therefore standard gradient-descent optimization is applicable. The advantage is better performance when compared to global optimization and also the smaller probe state set. Another advantage is that the global optimum is now unique, and the method unambiguously finds the true set $\{\hat{\pi}_{j}\}$ even when there is an unknown unitary operation between preparation and tomographic measurements. In Section S5 of the Supplementary material, we show a similar type of multiplicative error, where we optimize six parameters that describe the dependence of the optical phase on applied voltage in on-chip photonic circuits [27, 32, 33, 34].
Figure 2: Experimental demonstration for a polarization-encoded photonic qubit. Panels (a) and (b) show the purity and fidelity of reconstructed test states before (blue) and after (orange) application of measurement-device agnostic calibration. The fidelity of reconstructed states is calculated with respect to theoretical test states. The theoretical test states are depicted as empty black circles on the projected Bloch sphere in panels (c, d). Panel (c) depicts the reference case – the method was not used, and we assumed perfect waveplates in preparation and tomographic projections. In contrast, panel (d) shows the improved precision of preparation and measurement when we use the method and take into account the retardance deviation. The purity modulation is reduced and the test states match closely the theoretical expectations. In both panels, the reconstructed states are represented on the Bloch sphere using Hammer projection, and their purity is color-coded.
Experiment: polarization state tomography - We experimentally demonstrate the method in the waveplate-based polarization state tomography scenario, an example of a situation with multiplicative errors. Quarter-wave and half-wave plates transform the initial horizontal polarization of the photons, preparing the probe states with a uniform purity. The angular positions of the waveplates determine the resulting prepared state.
The tomography consists of projective measurements. First, the analyzed state undergoes unitary evolution provided by a half- and quarter-wave plate and is subsequently projected using a horizontally oriented linear polarizer. The action of the waveplate could be seen as the rotation of the Bloch sphere around axis $(\sin(2\alpha),0,\cos(2\alpha))$ with rotation angle $\Gamma$, where $\alpha$ determines the angular position of the waveplate and $\Gamma$ its retardance. The retardance usually slightly deviates from its nominal value. These deviations manifest as under- or over-rotation of the Bloch sphere. We parameterize the measurement operator by precisely controlled waveplate angular positions $x$, $y$, and half- and quarter-wave plate retardance deviations $\delta$ and $\epsilon$, respectively. These deviations are initially unknown to us, and we will use the presented method to estimate their values.
To demonstrate the tolerance of the method to a unitary operation between preparation and tomographic measurement, we add a quarter-wave plate with its fast axis oriented at -12.5 deg relative to the horizontal direction. We use eight probe states and perform the Pauli tomography for each. Initially, we assume zero deviations, $\delta=\epsilon=0$, and observe the purity modulation $\Delta P=0.06$. Then, we vary parameters $\delta,\epsilon$ and update the measurement operators employed in the reconstruction to minimize the purity modulation. The optimal $\delta$ and $\epsilon$ are close to the true values because the reconstruction artifact is significantly reduced to $\Delta P=0.01$. We found the optimal parameters $\delta=5.5$ deg, $\epsilon=-1.5$ deg.
We then reverse the process to determine retardance deviations $\tilde{\delta},\tilde{\epsilon}$ of the waveplates used in state preparation. We prepare the six eigenstates of Pauli operators, project each onto eight states corresponding to the original probe state set, and treat this data as before. We minimize the apparent purity modulation to find the optimal parameters $\tilde{\delta}=4.5$ deg and $\tilde{\epsilon}=-3.6$ deg, reducing the purity modulation from 0.07 to 0.01.
We verified the accuracy of the determined parameters using the following steps. We removed the central quarter-wave plate and then prepared an ensemble of test states, depicted as empty circles in Figure 2. As a reference, we first assumed perfect waveplates. The minimum fidelity of the reconstructed state to the target states was $F_{\min}=0.967$, and the average fidelity was $\bar{F}=0.992$. It indicates imprecise preparation. The purity modulation, $\bar{P}=0.990$ and $P_{\min}=0.938$, indicates reconstruction artifacts. When we used our knowledge of $\delta$ and $\epsilon$ in the reconstruction, the average purity increased to $\bar{P}=0.995$ and the minimum purity increased to $0.982$. The information gained in the calibration clearly mitigated the reconstruction artifact.
As a next step, we use this information to improve state preparation and measurement control. We calculated the new angular position of waveplates, considering the discovered retardance deviations, to implement the improved preparation and projection and repeated the experiment, obtaining $F_{\min}=0.991$, $\bar{F}=0.997$, $\bar{P}=0.995$, and $P_{\min}=0.983$. The results shown in Figure 2 indicate improved quality of state preparation and tomographic measurements. The small remaining artifacts are present due to the limited precision of the learned retardance deviations and also the repeatability of the waveplate rotation. Further, in Section S4 of the Supplementary material, we show the use of the presented method for classical rotating waveplate polarimetry [35].
Conclusion - The mismatch between actual measurement operators and their theoretical counterparts assumed in reconstructing quantum states or devices introduces reconstruction artifacts. We showed how to leverage these artifacts to reveal the actual measurement operators and mitigate systematic errors in the tomography and state preparation. We experimentally applied this method to Pauli tomography of polarization-encoded photonic qubits and numerically analyzed several other measurement schemes. The proposed method makes the tomography independent of the measurement device. The main advantage is that the method does not require perfect knowledge or control of the probe states, which is virtually impossible to obtain. The method can calibrate already existing tomographic apparatus without the need for individual characterization of its constituent components. This is particularly useful in the case where the experimental setup is monolithic and individual components cannot be calibrated individually, e.g. in the case of integrated circuits. One can also calibrate the state preparation with the reversed version of the method. Furthermore, the presented approach applies locally on individual parties of a larger multiparty system; i.e. it is fully scalable. In summary, we developed the framework for accurate and scalable calibration of quantum measurement devices and also quantum state preparation, which is critical in any applications of quantum science and technology.
The application of the reported measurement-device agnostic quantum measurement goes beyond the full quantum tomography. The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches such as Monte-Carlo sampling [36, 37], permutationally invariant tomography [38], compressed sensing [39, 40], etc. The method can also improve the accuracy of building and characterization of the photonic and quantum circuits in complex media [41, 42]. Finally, the reported approach also applies to experimental realizations of entanglement certification and quantification [23, 24, 22, 43, 25].
The supporting data for this article are available from the Zenodo repository [44].
Acknowledgments - We acknowledge the support of the Czech Science Foundation under grant No. 21-18545S. M. B. acknowledges the support of Palacký University under grant No. IGA-PrF-2024-008. We acknowledge the use of cluster computing resources provided by the Department of Optics, Palacký University Olomouc. We thank J. Provazník for maintaining the cluster and providing support.

### v2

@@
 The supporting data for this article are available from the Zenodo repository [44].
+Acknowledgements.
 Acknowledgments - We acknowledge the support of the Czech Science Foundation under grant No.
@@
 Provazník for maintaining the cluster and providing support.
+Supplemental Material: Measurement-device agnostic quantum tomography
+I NOMAD setting in the numerical test
+When we numerically investigated the case of general additive errors, we set the NOMAD [1] parameters as listed in Tab.
+1.
+Please refer to the NOMAD 4 manual [2] for the meaning of the parameters.
+The optimized parameters were bounded to the interval (-29, 29) deg, i.e., roughly three standard deviations of the true parameter distribution.
+The seed was pseudo-randomly chosen at the time of optimization.
+The maximum size of the evaluation block was set as twice the problem dimension to allow optimal parallelization of the initial Latin hypercube search.
+We used the nomadlad [3] package to use NOMAD from Python.
+The necessary configuration was to set NOMAD to use multiple initial optimization points instead of just one by not specifying parameter X0 and setting the Latin hypercube search parameter (LH SEARCH) sufficiently high.
+parameter value
+BB OUTPUT TYPE OBJ
+DIMENSION 12
+BB MAX BLOCK SIZE 24
+MAX BB EVAL 100 000
+SEED random integer
+DIRECTION TYPE ORTHO 2N
+LH SEARCH 1100 1100
+LOWER BOUND (-0.5, …, -0.5)
+UPPER BOUND (0.5, …, 0.5)
+Table 1: NOMAD (v4.1.0) settings used for the global optimization. Some parameters are not listed because they are not critical for this application.
+II Artifact mitigation of multiplicative errors
+Here, we explore the sampling strategies in the case of multiplicative errors.
+Since the artifact is state-dependent, the probe sampling strategy is important for minimizing the reconstruction artifacts.
+As in the main text, all states are measured in projections onto eigenstates of Pauli operators.
+The projections are implemented by rotating the measured state using the following unitary operation $\hat{R}_{y}(\theta_{j})\hat{R}_{z}(\phi_{j})$ and subsequent projection of the resulting state onto state $|0\rangle$.
+The rotation angles $\theta_{j},\phi_{j}$ determine the selected projection and are listed in Tab 2.
+We test a few sampling schemes and plot the cost function near the ground truth ($\delta^{\prime}=0.02$, $\epsilon^{\prime}=-0.04$) for each.
+The cost function is the peak-to-valley value of reconstructed probe state purities, $\Delta P=P_{\max}-P_{\min}$.
+The probe sets and cost functions are depicted in Figure S1.
+As probe states set, we selected eigenstates of Pauli operators (panels a,b), eight states in (panels c, d), icosahedron (panels e, f), and then a uniform sampling of Bloch sphere angular coordinates ($\theta,\phi$) with 14 states (panels g, h) and 22 states (panels i, j) and finally NASA HEALPIX [4, 5] with 108 probe states (panels k, l).
+$j$ 1 2 3 4 5 6
+$\theta_{j}$ ($y$-axis) 0 $\pi$ $\pi/2$ $\pi/2$ $\pi/2$ $\pi/2$
+$\phi_{j}$ ($z$-axis) 0 0 $\pi$ 0 $\pi/2$ $23\pi/$ 2
+Table 2: Rotation angles in radians for Pauli tomography.
+The analyzed state is first rotated by $\phi_{j}$ around the $z$-axis and $\theta_{j}$ around the $y$-axis and subsequently projected to state $|0\rangle$.
+The sampling is sufficient when the cost function possesses a unique global minimum.
+We have run additional simulations, varied the ground truth, and observed the shape of the cost function.
+We tested each sampling scheme from Fig.
+S1.
+We sampled the cost function near the varying ground truth values of $\delta^{\prime}$ and $\epsilon^{\prime}$ and plotted its value as color maps, depicted in Fig.
+S1.
+The first suitable sampling scheme appears for eight probe states.
+Interestingly, for 14 probe states, there is a rift in the cost function landscape that causes ambiguity in the $\delta,\epsilon$ determination.
+The remaining sampling schemes exhibit a well-defined global minimum and therefore are suitable.
+The cost function landscapes in Fig.
+S1 are plotted for a single choice of $(\delta,\epsilon)$.
+To better characterize the sampling strategies, we repeated the same simulations for multiple ground truths.
+Figures S2-S6 show the simulated cost function for 8, 12, 17, 22, and 108 probe states.
+The ground truth is always plotted in the panel center.
+This additional characterization confirms the result – 8 probe states are enough and there is indeed anomalous behavior for 14 probe states.
+For 22 states, we sometimes see a local minimum, such as for ground truth parameters $\delta^{\prime}=0.15$, $\epsilon^{\prime}=0$.
+Due to this, the optimization has to be done carefully, ideally with several starting points, to find the global minimum of the cost function.
+In the case of 108 probe states, the cost function always had a single minimum.
+The most interesting outcome of this analysis is that the success of the sampling strategy is not determined by the sheer number of probe states but rather by their suitable placement on the Bloch sphere.
+Generally, adding more probe states makes the error parameters determination more robust because we increase the obtained information, however, at the cost of longer measurement time.
+The exact cost function shape, sampling strategies, and shapes of the purity modulation depend on the imperfection type, its parameterization, and how we switch the measurement operators in the tomography.
+Figure S1: Dependence of simulated artifact severity, represented as purity modulation amplitude depicted as 2D color maps in panels (a, c, …, k) on the selection of probe states set, depicted as points on projected Bloch sphere in panels (b, d, …, l).
+The figure shows the case for 6 (a,b), 8 (c,d), 12 (e,f), 14 (g,h), 22 (i,j), and 108 (k,l) probe states.
+The purity modulation amplitude is always plotted around the ground truth $\delta=0.02$, $\delta=-0.04$.
+Figure S2: The panels show color-coded purity modulation as a function of assumed error parameters ($\delta,\epsilon$) observed in the set of eight reconstructed probe states.
+The purity modulation is sampled near the true value ($\delta^{\prime},\epsilon^{\prime}$) for each simulated ground truth.
+This true value is always plotted in the center of the panel.
+All panels share the color bar.
+To visually emphasize the presence of a single minimum, we use logarithmic normalization of color coding.
+Figure S3: 12 probe states. The figure is arranged as Fig. S2.
+Figure S4: 14 probe states. The figure is arranged as Fig. S2.
+Figure S5: 22 probe states. The figure is arranged as Fig. S2.
+Figure S6: 108 probe states. The figure is arranged as Fig. S2.
+III Experimental setup details
+We used a continuous-wave 810 nm fiber-coupled laser diode of 1 mW power.
+The laser diode was thermally stabilized using Peltier elements.
+The laser beam was decoupled into free space using an 11-mm aspheric lens and horizontally polarized using a Glan-Taylor prism.
+The angular positions of quarter-wave and half-wave plates controlled the resulting polarization of the beam.
+The following operator describes the action of a waveplate:
+$\hat{W}_{\alpha,\gamma}=|\alpha\rangle\langle\alpha|+e^{-i\Gamma}|\alpha_{\bot}\rangle\langle\alpha_{\bot}|,$ (S1)
+where $\alpha$ and $\Gamma$ are its angular position and retardance, and $|\alpha\rangle=\cos{\alpha}|0\rangle+\sin{\alpha}|1\rangle$ is a linearly polarized state, and $\alpha_{\bot}=\alpha+\pi/2$.
+To prepare a state located at colatitude $\theta$ and longitude $\phi$ on the Bloch sphere, we rotate the half- and quarter-wave plate to angles $x$ and $y$ given by the relations
+$y=-\frac{1}{2}\arcsin\left(\sin\theta\sin\phi\right),$ (S2)
+$x=\frac{1}{4}\arctan\left(\tan\theta\cos\phi\right)+\frac{1}{2}y+c(\theta),$ (S3)
+where $c(\theta)=\pi/4$ if $\theta>\pi/2$ and $c(\theta)=0$ otherwise.
+This relation holds for wave plates that have perfect retardance.
+The deviation from the ideal retardance will introduce errors in the preparation.
+We assumed the perfect retardance in the first step of the experiment as our method does not require exact knowledge of probe states.
+Then we inserted a rotated quarter-wave plate that rotated the prepared state to demonstrate the robustness of our method to unitary operations in between the preparation and projection stages.
+The states were analyzed using half-waveplate and quarter-wave and subsequent projection onto horizontal and vertical polarization using a calcite beam-displacer.
+The waveplate settings for the Pauli tomography are listed in Tab.
+3.
+All waveplates were mounted into a precision motorized rotation stage that provided bidirectional repeatability of ±$\pm 30$ mdeg.
+The optical signals were measured by reading photo-currents from a pair of reverse-biased photodiodes.
+We normalized the readings using the sum of two signals to eliminate laser power fluctuation.
+The same results are obtained for a single-photon signal using two single-photon detectors.
+The calibration step revealed the retardance deviation of the preparation and projection wave plates.
+To reconstruct the quantum states, we could directly use this knowledge to update our measurement operators.
+However, we take one extra step which allows us to prepare the states with greater precision and control the measurement operators precisely.
+We took these deviations into account in the verification step and calculated new angular positions of wave plates for tomographic projections and state preparation.
+Since Eq.
+S3 does not hold anymore, we used numerical optimization to find the desired wave plate settings.
+To prepare a state $|\psi\rangle$, we find the wave plate angles $x,y$ using the following maximization:
+$x,y=\mathrm{argmax}_{x,y}|\langle\psi|\hat{W}(x,\pi+\delta)\hat{W}(y,\pi/2+\epsilon)|0\rangle|^{2}.$ (S4)
+Similarly, to find the settings to project a state onto state $|\pi\rangle$ we perform this maximization
+$x,y=\mathrm{argmax}_{x,y}|\langle\langle 0|\hat{W}(x,\pi/2+\epsilon)\hat{W}(y,\pi+\delta)|\pi\rangle|^{2}.$ (S5)
+state $|H\rangle$ $|V\rangle$ $\frac{|H\rangle+|V\rangle}{\sqrt{2}}$ $\frac{|H\rangle-|V\rangle}{\sqrt{2}}$ $\frac{|H\rangle+i|V\rangle}{\sqrt{2}}$ $\frac{|H\rangle-i|V\rangle}{\sqrt{2}}$
+$j$ 1 2 3 4 5 6
+$x_{j}$ (HWP) 0 $\pi/4$ $\pi/8$ $-\pi/8$ $-\pi/8$ $\pi/8$
+$y_{j}$ (QWP) 0 0 0 0 $-\pi/4$ $\pi/4$
+$x^{\prime}_{j}$ (HWP) 0 $\pi/4$ $\pi/8$ $-\pi/8$ $\pi/8$ $-\pi/8$
+$y^{\prime}_{j}$ (QWP) 0 0 0 0 $\pi/4$ $-\pi/4$
+Table 3: Waveplate angular position for Pauli tomography.
+The waveplates are placed as follows in the direction of the light propagation: quarter-wave plate and half-wave plate in the preparation, and half-wave plate and quarter-wave plate in the measurement stage.
+Primed symbols correspond to the state preparation, used in the reversed version of the method.
+IV Rotating waveplate polarimetry
+Consider the classical rotating quarter-wave plate polarimeter [6], depicted in Fig.
+S7(a), with the quarter-wave plate having the true retardance deviated by $\delta^{\prime}=5\deg$ from its nominal value.
+Moreover, consider pure probe states.
+The waveplate rotates continuously and effectively modulates the optical signal detected after the polarizer.
+Instead of likelihood maximization, we now use classical Fourier analysis to reconstruct the polarization state.
+We chose the Fourier analysis to show that our method is not limited to maximal likelihood reconstruction.
+We used eight probe states to detect the reconstruction artifact, and instead of purity modulation, we equivalently measured the severity with a modulation amplitude of the degree of polarization.
+First, we treat the data with the ideal retardance ($\delta=0$) assumption.
+We see that the states from the ensemble suffer from artifacts.
+Here, we look at the intensity-normalized Stokes vector length $D=\frac{\sqrt{\sum\limits_{i=x,y,z}S_{i}}}{S_{0}}$, which is proportional to purity $P=\frac{1}{2}\left(1+D^{2}\right)$.
+Near state |$\ket{H}$, at the north pole of the Bloch sphere, the purity even exceeds 1 due to the reconstruction method.
+The modulation of $D$ is depicted in Fig.
+S7(c).
+This behavior is clearly a reconstruction artifact.
+Exceeding the unit purity is unsurprising because Fourier analysis does not guarantee physically sound results.
+We now vary $\delta$ assumed in reconstruction and observe the peak-to-peak value $\Delta D$ in the probe set.
+The dependence is depicted in Fig.
+S7(b).
+We see that the minimum lies at the true $\delta^{\prime}$.
+In panel (d), we show that the artifact disappears when the retardance deviation is taken into account.
+Figure S7: Mitigation of artifact in rotating-waveplate polarimetry.
+(a) Measurement apparatus.
+A quarter-wave plate (purple) with retardance deviation $\delta^{\prime}=5$ deg is rotating in front of the linear polarizer.
+The photodiode signal is recorded and analyzed.
+(b) The minimal $\Delta D$ lies at the true retardance deviation $\delta^{\prime}$.
+(c) Initial mismatch results in state-dependent purity artifact.
+We plot the reconstructed probe states on the projected Bloch sphere and color-code their reconstructed stokes length $D=\sqrt{S^{2}_{x}+S^{2}_{y}+S^{2}_{z}}/S_{0}$.
+(d) When we considered the true retardance, the artifact vanished.
+Black circles depict the ideal testing states.
+V Simulated experiment with path-encoded qubits on a photonic chip
+As another example, we illustrate the measurement-device agnostic calibration for tomography of photonic on-chip path qubits [7, 8, 9, 10].
+The qubit is encoded into the photon propagation in one or another waveguide (states $\ket{\uparrow}$ or $\ket{\downarrow}$), as depicted in Figure S8(a).
+The local heating of one waveguide controls the phase between the two paths.
+The heating is typically achieved by an electrical current flowing through a small heater element.
+A Mach-Zehnder interferometer serves as a variable coupler, and its interferometric phase controls the coupling ratio.
+Detection of photons in a single optical path implements a projective measurement of the path qubit.
+The circuit in Figure S8(a) implements a projection onto state |$\ket{\pi}$, whose colatitude and longitude on the Bloch sphere are controlled by optical phases.
+The relation between optical phases $\varphi_{1,2}$ and voltages $V_{1,2}$ applied on the heating elements is approximated with polynomials [8]
+$\varphi_{1}(V_{1})=c_{1}+c_{2}V_{1}+c_{3}V_{1}^{2},$ (S6)
+$\varphi_{2}(V_{2})=c_{4}+c_{5}V_{2}+c_{6}V_{2}^{2}.$ (S7)
+Figure S8: Application of the method for on-chip photonic path qubit.
+(a) Projection apparatus for a path qubit.
+Applied voltage $V_{1,2}$ on heater elements (orange) causes thermal-based phase shift in the waveguide interferometers.
+These shifts determine the measurement projector.
+The photons in the upper output path are detected with a single-photon detector (SPAD).
+(b) 1D-section of 6D function $P_{\min}(c_{1},\dots,c_{6})$.
+The value of $P_{\min}$ at the initial assumption of $c_{j}$ is 0.955.
+Maximum lies at the true value $c^{\prime}_{6}$.
+(c) Initial mismatch results in state-dependent purity artifact.
+We plot probe states on the Bloch sphere and color-code their reconstructed purity.
+(d) When we consider the true retardance, the artifact vanishes.
+Parameters of the simulation are: $c^{\prime}_{1}=-1.887$, $c^{\prime}_{2}=0.105$ V-1, $c^{\prime}_{3}=0.05$ V-2, $c^{\prime}_{4}=-1.805$, $c^{\prime}_{5}=0.115$ V-1, and $c^{\prime}_{6}=0.05$ V-2.
+Initially assumed parameters were: $c_{1}=-1.750$, $c_{2}=0.100$ V-1, $c_{3}=0.050$ V-2, $c_{4}=-1.800$, $c_{5}=0.1105$ V-1, and $c_{6}=0.050$ V-2
+Here, the voltages $V_{1,2}$ are controlled perfectly, but parameters $c_{j}$ are known only approximately.
+The other imperfections are neglected.
+We have increased the number of probe states to 22 to compensate for the higher number of optimization parameters.
+The simulated experiment, depicted in Fig.
+S8(c), shows the measurement operator mismatch artifact.
+We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j})$ where the systematic error is mitigated.
+Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c^{\prime}_{6}$ (indicated by the dashed line).
+Taking the true values $c^{\prime}_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).

### v4

@@
 It consequently allows the mitigation of systematic errors in quantum measurement and state preparation.
-Introduction - The measurements play a central role in quantum science and technology [1].
+I Introduction
+The measurements play a central role in quantum science and technology [1].
 However, the implemented measurements often differ from the ideal ones, leading to systematic errors.
@@
 The framework suggests the existence of a self-calibrating state class, which could reveal, when measured, the initially unknown information about the states and the measurement device.
-The cases of photon-number-resolving measurement and quantum homodyne tomography have been discussed.
 Another approach demonstrated for polarization states of light infers a single unknown parameter of the measurement device using maximum likelihood estimation [12].
+There are also schemes that use simultaneous likelihood maximization over both states and devices [13] or states and measurements [14].
 Finally, assuming the sparsity of the parameter vector and the low rank of the measured density matrix, the dependence on the measurement device can be relaxed [15].
-In this Letter, we show that an ensemble of single-qubit states of equal purity distributed quasi-uniformly on the Bloch sphere forms a class of self-calibrating states.
+In this work, we show that an ensemble of single-qubit states of equal purity distributed quasi-uniformly on the Bloch sphere forms a class of self-calibrating states.
 This class is convenient in experiments where states are naturally prepared with quasi-equal purity, e.g., in photonics and trapped ions.
@@
 We show a novel approach to simultaneously calibrating quantum measurements and state preparation.
-To this end we exploit artifacts in quantum state reconstruction performed on the self-calibrating probe states.
+To this end, we exploit artifacts in quantum state reconstruction performed on the self-calibrating probe states.
 In the multi-qubit setting, we can use the method in a fully scalable way to mitigate systematic errors on each local qubit measurement, effectively mitigating errors in the whole system.
@@
 Figure 1: Purity modulation due to the measurement operator mismatch.
-States are represented as points on the Bloch sphere plotted using Hammer projection, and the purity of their reconstruction is color-coded.
+(a-c) States are represented as points on the Bloch sphere plotted using Hammer projection, and the purity of their reconstruction is color-coded.
 Although all probe states were prepared with the same purity, the reconstruction using quantum tomography with mismatched measurement operators yields purity modulation.
-Measurement operator mismatch artifact - In quantum state tomography, many copies of an initially unknown state, described by density operator $\hat{\rho}$, are subject to a series of measurements described by measurement operators $\hat{\pi}_{j}$.
+The purity modulation scales with the magnitude of error parameters.
+Panel (b) shows the purity modulation for randomly selected error parameters described in the main text.
+Panels (a) and (c) show how the purity modulation, if we scale the error vector by a factor of $\lambda=0.25$ and $\lambda=4$, respectively.
+The black curve in panel (d) depicts the scaling trend, and red dots on the curve mark the selected values of $\lambda$ plotted in panels (a-c).
+II Measurement operator mismatch artifact
+In quantum state tomography, many copies of an initially unknown state, described by density operator $\hat{\rho}$, are subject to a series of measurements described by measurement operators $\hat{\pi}_{j}$.
 These measurement operators should be tomographically complete, and their choice influences the performance of the tomography [18, 19, 20, 21].
@@
 The tomogram and the measurement operators are then input into an algorithm, which outputs the density operator.
-If the actual measurement operators $\hat{\pi}^{\prime}_{j}$ differ from those input into the reconstruction method, artifacts appear.
-We use the maximum-likelihood method [20], but in Section S4 of the Supplementary Material, we demonstrate that the effect is not limited to the particular reconstruction method.
-The disagreement between actual and assumed measurement operators causes inconsistency in the tomogram, which results in state-dependent purity loss.
+Here, we use the maximum-likelihood reconstruction algorithm [13, 22].
+If the actual measurement operators $\hat{\pi}^{\prime}_{j}$ differ from those put into the reconstruction method, artifacts appear.
 Let us demonstrate the behavior in the simplest scenario of pure probe single-qubit states and projective measurements onto the eigenstates of Pauli operators, called Pauli tomography.
@@
 $p_{j}=\mathrm{Tr}[\hat{\rho}\hat{\pi}_{j}].$ (2)
-The angles $\theta_{j}$ and $\phi_{j}$ for Pauli tomography are provided in Table S2 of the Supplemental Material. Assume that due to experimental imperfections, the actual angles
+The angles $\theta_{j}$ and $\phi_{j}$ for Pauli tomography are provided in Table S2 of the Supplemental Material. Assume that due to experimental imperfections, the actual angles are
 $\theta_{j}^{\prime}=\theta_{j}+\delta_{j},$ (3)
@@
 Then, we reconstruct them assuming the original unperturbed measurement operators $\{\hat{\pi}_{j}\}$.
-This mismatch causes the state-dependent purity modulation, shown in Figure 1.
+This mismatch causes the state-dependent purity modulation, shown in Figure 1(a-c).
+This example shows the apparent purity modulation for the case of an error vector $\vec{\delta}$, whose elements are drawn from a normal distribution with $\sigma=5$ deg.
 We quantify the purity modulation in the ensemble of reconstructed states as
@@
 where $P=\mathrm{Tr}\left(\hat{\rho}^{2}\right)$ is the purity of the reconstructed state.
-The modulation $\Delta P$ scales with the magnitude of parameters $\vec{\delta}$ and thus quantifies the severity of the measurement operator mismatch.
-The error vector $\vec{\delta}$ determines at which state the artifact manifests maximally.
-We will leverage these features to discover the vector $\delta$ and, consequently, the true measurement operators.
-Artifact mitigation - We minimize the artifact severity, i.e., purity modulation (5), by varying the measurement operators used to reconstruct the probe states.
+To show how purity modulation scales with the error vector magnitude, we scale it with a non-negative real factor $\lambda$, i.e.
+$\vec{\delta}\rightarrow\lambda\vec{\delta}$, and observe the $\Delta P$, as shown in Figure 1(d).
+We see that $\Delta P$ monotonously increases with $\lambda$.
+Consequently, $\Delta P$ effectively measures how far the assumed measurement operators are from the actual ones.
+Moreover, the shape of the purity landscape is determined by the error vector $\vec{\delta}$.
+We will leverage this dependence of apparent purity as a function of vector $\delta$ to find the true measurement operators.
+The method is not limited to the particular reconstruction method, as we illustrate in an example of rotating quarter-wave plate polarimetry with classical Fourier retrieval of Stokes parameters [23].
+The example is provided in Supplementary Material S4.
+In that case, the disagreement between actual and assumed measurement operators causes inconsistency in the tomogram, which results in apparent state-dependent modulation of the degree of polarization, where the degree of polarization can even exceed one.
+III Artifact mitigation
+We minimize the artifact severity, i.e., purity modulation (5), by varying the measurement operators used to reconstruct the probe states.
 After recording all tomograms of the probe states we used, we optimize our assumption of $\vec{\delta}$ which specifies the measurement operators used in the reconstruction algorithm.
@@
 The probe states should cover the Bloch sphere quasi-uniformly and possess uniform purity; otherwise, their full control is not required.
-This feature is advantageous in practice because state preparation is never perfect in real-world implementation.
+This feature is advantageous in practice because state preparation is never perfect in the experiment.
 It distinguishes our approach from the schemes that rely on perfect knowledge of the probe states, like quantum measurement tomography [8].
@@
 The resulting optimum $\vec{\delta}$ is not unique because any set of measurement operators $\{\hat{U}\hat{\pi}_{j}\hat{U}^{\dagger}\}$, where $\hat{U}$ is a fixed unitary operation, also minimizes the artifact.
-Each of these satisfactory sets corresponds to different parameters $\vec{\delta}$.
 These unitary-equivalent solutions could be interpreted as unitary transformations of probe states.
@@
 The sampling with too few probe states could lead to false error parameters $\vec{\delta}_{f}$.
-We used 30 quasi-uniformly distributed probe states.
-Such a choice provided a reliable determination of error parameters.
-We discuss the effects of the number of probe states in detail in the Supplementary material.
+We found out that 30 quasi-uniformly distributed probe states provide a reliable determination of error parameters in all tested cases.
+When we used fewer probe states, we encountered incorrect estimates.
+The incorrect estimate provides reduced apparent purity modulation in the set of probe states, but when it is used on other states, the purity modulation can be even larger than initially.
 To test the method numerically, we first randomly generate ground-truth error parameters $\vec{\delta}$.
@@
 For simplicity, we exclude statistical noise from our simulations.
-Parameters $\vec{\delta}^{\prime}$ assumed in the reconstruction are optimized to minimize the purity modulation $\Delta P$ in the set of reconstructed probe states $\{\hat{\rho}_{k}\}_{k=1,\dots,30}$, concluding the calibration step.
+Parameters $\vec{\delta}^{\prime}$ assumed in the reconstruction are optimized to minimize the purity modulation $\Delta P$ in the set of reconstructed probe states $\{\hat{\rho}_{k}\}$ for $k=1,\dots,30$, concluding the calibration step.
 To verify the correctness of the optimization result, we use the ground truth $\vec{\delta}$ again to generate tomograms of pure test states $\hat{\xi}=|\xi_{k}\rangle\langle\xi_{k}|$, which quasi-uniformly covers the Bloch sphere in 108 points.
@@
 We apply $\hat{W}$ to all reconstructed test states $\hat{\xi}_{k}$ to eliminate the unitary-related ambiguity.
-Then, we check their fidelity with the expected ideal test state $F_{k}=\langle\xi_{k}|\hat{\rho}_{k}|\xi_{k}\rangle$.
+Then, we check their fidelity with the expected ideal test state, $F_{k}=\mathrm{Tr}[\sqrt{\sqrt{\hat{\xi}_{k}}\hat{\rho}_{k}\sqrt{\hat{\xi}_{k}}}]$.
 We characterize the test states ensemble with the purity modulation and their lowest fidelity $\min\limits_{k}F_{k}$ to provide the worst-case estimate.
@@
 When we compare the optimized results to the reference, i.e.
-the states reconstructed with the assumption of null error vector, the purity modulation decreases from $\Delta P=(8\begin{subarray}{c}+3\\ -3\end{subarray}))\cdot 10^{-2}$ to $\Delta P=(1.0\begin{subarray}{c}+0.8\\ -1.0\end{subarray})\cdot 10^{-3}$.
-The uncertainty interval spans from 0.158 to 0.842-quantile in correspondence to $\pm$ one standard deviation and is used due to the skewed distribution of the results.
+the states reconstructed with the assumption of null error vector, the average purity modulation decreases from $\Delta P=(8\begin{subarray}{c}+3\\ -3\end{subarray}))\cdot 10^{-2}$ to $\Delta P=(1.0\begin{subarray}{c}+0.8\\ -1.0\end{subarray})\cdot 10^{-3}$.
+The uncertainty interval spans from 0.158 to 0.842-quantile in correspondence to one standard deviation and is used due to the skewed distribution of the results.
 Mainly, the infidelity decreased from $(3\begin{subarray}{c}+1\\ -2\end{subarray})\cdot 10^{-2}$ to $(3\begin{subarray}{c}+5\\ -3\end{subarray})\cdot 10^{-4}$, which represents two orders of magnitude improvement in accuracy.
@@
 It is important to stress that the pure probe states and projective (pure) measurements were chosen here for the simplicity of introducing the measurement-device agnostic framework.
-However, the method is fully applicable in real cases of nonunity purity (as in the following experimental demonstration) and generic positive-operator-valued measurements.
+However, the method is fully applicable in real cases of generic positive-operator-valued measurements and nonunity purity, as indicated in the following experimental demonstration.
 In the multi-qubit scenario, the tomography is usually done using local measurements.
@@
 The artifact severity scales worst as $P_{\min,n}=\prod\limits_{i=1}^{n}P_{\min,i}$ for $n$-qubit state tomography, where $P_{\min,i}$ is the corresponding minimal purity observed at the single qubit.
-If the input state |$\ket{\psi_{i}}$ at the $i$-th qubit suffers from the artifact the worst, then product state $\bigotimes\limits_{i}\ket{\psi_{i}}$ is the $n$-qubit state with the greatest manifestation of the artifact.
+If the input state |$\ket{\psi_{i}}$ at the $i$-th qubit suffers from the artifact the most, then product state $\bigotimes\limits_{i}\ket{\psi_{i}}$ is the $n$-qubit state with the greatest manifestation of the artifact.
 In such a case, the single-qubit improvements multiply and lead to significantly better overall results in multi-qubit scenarios.
-State preparation - The state preparation typically suffers from imperfections too.
+IV State preparation
+The state preparation typically suffers from imperfections too.
 Here, we use the symmetry in the Born rule to formally exchange state preparation and measurement to utilize the previously introduced method to correct state preparation.
@@
 Either way, we gain information about $\hat{U}_{j}$, increasing the accuracy of state preparation.
-Special case: multiplicative error - An important subclass of the problem is the case of multiplicative measurement operator error
+V Special case: multiplicative error
+An important subclass of the problem is the case of multiplicative measurement operator error, where
 $\theta_{j}^{\prime}=(1+\delta)\theta_{j},$ (6)
@@
 Another advantage is that the global optimum is now unique, and the method unambiguously finds the true set $\{\hat{\pi}_{j}\}$ even when there is an unknown unitary operation between preparation and tomographic measurements.
-In Section S5 of the Supplementary material, we show a similar type of multiplicative error, where we optimize six parameters that describe the dependence of the optical phase on applied voltage in on-chip photonic circuits [7, 8, 9, 10].
 Figure 2: Experimental demonstration for a polarization-encoded photonic qubit.
-Panels (a) and (b) show the purity and fidelity of reconstructed test states before (blue) and after (orange) application of measurement-device agnostic calibration.
+Panel (a) shows the experimental scheme: LD - laser diode, HWP (QWP) - half-(quarter)-wave plate, GT - Glan-Taylor polarizer, BD - calcite beam displacer, PD - photodiode.
+Panels (b) and (c) show the purity and fidelity of reconstructed test states before (blue) and after (orange) application of measurement-device agnostic calibration.
 The fidelity of reconstructed states is calculated with respect to theoretical test states.
-The theoretical test states are depicted as empty black circles on the projected Bloch sphere in panels (c, d).
-Panel (c) depicts the reference case – the method was not used, and we assumed perfect waveplates in preparation and tomographic projections.
-In contrast, panel (d) shows the improved precision of preparation and measurement when we use the method and take into account the retardance deviation.
+The theoretical test states are depicted as empty black circles on the projected Bloch sphere in panels (d, e).
+Panel (d) depicts the reference case – the method was not used, and we assumed perfect waveplates in preparation and tomographic projections.
+In contrast, panel (e) shows the improved precision of preparation and measurement when we use the method and take into account the retardance deviation.
 The purity modulation is reduced and the test states match closely the theoretical expectations.
 In both panels, the reconstructed states are represented on the Bloch sphere using Hammer projection, and their purity is color-coded.
-Experiment: polarization state tomography - We experimentally demonstrate the method in the waveplate-based polarization state tomography scenario, an example of a situation with multiplicative errors.
-Quarter-wave and half-wave plates transform the initial horizontal polarization of the photons, preparing the probe states with a uniform purity.
-The angular positions of the waveplates determine the resulting prepared state.
-The tomography consists of projective measurements.
-First, the analyzed state undergoes unitary evolution provided by a half- and quarter-wave plate and is subsequently projected using a horizontally oriented linear polarizer.
+VI Experiment: polarization state tomography
+We experimentally demonstrate the method in the waveplate-based polarization state tomography scenario, an example of a situation with multiplicative errors.
+At the end of the process, we will improve our knowledge of the waveplate retardance and consequently mitigate the reconstruction artifact.
+The experimental setup is depicted in Fig.
+1(a).
+We used a continuous-wave 810 nm thermally-stabilized fiber-coupled laser diode (LD).
+The laser beam was decoupled into free space using an 11-mm aspheric lens and horizontally polarized using a Glan-Taylor prism (GT).
+The waveplates HWP1 and QWP1 effectively control the prepared polarization state.
+The following operator describes the action of a waveplate:
+$\hat{W}_{\alpha,\gamma}=|\alpha\rangle\langle\alpha|+e^{-i\Gamma}|\alpha_{\bot}\rangle\langle\alpha_{\bot}|,$ (8)
+where $\alpha$ and $\Gamma$ are its angular position and retardance, and $|\alpha\rangle=\cos{\alpha}|0\rangle+\sin{\alpha}|1\rangle$ is a linearly polarized state, and $\alpha_{\bot}=\alpha+\pi/2$.
+To prepare a state located at colatitude $\theta$ and longitude $\phi$ on the Bloch sphere, we rotate the half- and quarter-wave plate to angles $x^{\prime}$ and $y^{\prime}$ given by the relations
+$y^{\prime}=-\frac{1}{2}\arcsin\left(\sin\theta\sin\phi\right),$ (9)
+$x^{\prime}=\frac{1}{4}\arctan\left(\tan\theta\cos\phi\right)+\frac{1}{2}y+c(\theta),$ (10)
+where $c(\theta)=\pi/4$ if $\theta>\pi/2$ and $c(\theta)=0$ otherwise.
+This relation holds for wave plates that have perfect retardance.
+The deviation from the ideal retardance will introduce errors in the preparation.
+We assumed the perfect retardance in the first step of the experiment as our method does not require exact knowledge of probe states.
+To demonstrate the tolerance of the method to a unitary operation between preparation and tomographic measurement, we add a quarter-wave plate (QWPX) with its fast axis oriented at -12.5 deg relative to the horizontal direction.
+Because this unitary operation can be in principle unknown, we restrict ourselves from using any information about QWPX in the calibration.
+The states were analyzed using half-waveplate and quarter-wave (HWP2, QWP2) and subsequent projection onto horizontal and vertical polarization using a calcite beam-displacer (BD).
+The waveplate settings for the Pauli tomography are listed in Tab.
+S2 of the Supplementary Material.
+All waveplates were mounted into a precision motorized rotation stage that provided bidirectional repeatability of ±$\pm 30$ mdeg.
+We used reverse-biased PIN photodiodes (PD) to measure optical intensity at both outputs of the beam-displacer.
+We normalized the readings using the sum of two signals to eliminate laser power fluctuation.
 The action of the waveplate could be seen as the rotation of the Bloch sphere around axis $(\sin(2\alpha),0,\cos(2\alpha))$ with rotation angle $\Gamma$, where $\alpha$ determines the angular position of the waveplate and $\Gamma$ its retardance.
@@
 These deviations are initially unknown to us, and we will use the presented method to estimate their values.
-To demonstrate the tolerance of the method to a unitary operation between preparation and tomographic measurement, we add a quarter-wave plate with its fast axis oriented at -12.5 deg relative to the horizontal direction.
 We use eight probe states and perform the Pauli tomography for each.
-Initially, we assume zero deviations, $\delta=\epsilon=0$, and observe the purity modulation $\Delta P=0.06$.
-Then, we vary parameters $\delta,\epsilon$ and update the measurement operators employed in the reconstruction to minimize the purity modulation.
-The optimal $\delta$ and $\epsilon$ are close to the true values because the reconstruction artifact is significantly reduced to $\Delta P=0.01$.
-We found the optimal parameters $\delta=5.5$ deg, $\epsilon=-1.5$ deg.
+Initially, we assume zero deviations, $\delta=\epsilon=0$, and observe the purity modulation $\Delta P=0.0572(3)$.
+The number in parenthesis is one standard deviation at the last significant digit, e.g.
+$0.0572(3)=(572\pm 3)\cdot 10^{-4}$.
+The uncertainties were estimated using bootstrapping with the assumption of normally distributed photo-current with 0.1% relative standard deviation, which is the most dominant term and is related to the ammeter precision.
+The shot noise is negligible at the power level of our laser, compared to the technical noise.
+Then, we vary parameters $\delta$ and $\epsilon$ and update the measurement operators employed in the reconstruction to minimize the purity modulation.
+The optimal $\delta$ and $\epsilon$ are close to the true values because the reconstruction artifact is significantly reduced to $\Delta P=0.0140(2)$.
+We found the optimal parameters $\delta=4.5(2)$ deg, $\epsilon=-1.30(4)$ deg.
 We then reverse the process to determine retardance deviations $\tilde{\delta},\tilde{\epsilon}$ of the waveplates used in state preparation.
 We prepare the six eigenstates of Pauli operators, project each onto eight states corresponding to the original probe state set, and treat this data as before.
-We minimize the apparent purity modulation to find the optimal parameters $\tilde{\delta}=4.5$ deg and $\tilde{\epsilon}=-3.6$ deg, reducing the purity modulation from 0.07 to 0.01.
-We verified the accuracy of the determined parameters using the following steps.
-We removed the central quarter-wave plate and then prepared an ensemble of test states, depicted as empty circles in Figure 2.
-As a reference, we first assumed perfect waveplates.
-The minimum fidelity of the reconstructed state to the target states was $F_{\min}=0.967$, and the average fidelity was $\bar{F}=0.992$.
-It indicates imprecise preparation.
-The purity modulation, $\bar{P}=0.990$ and $P_{\min}=0.938$, indicates reconstruction artifacts.
-When we used our knowledge of $\delta$ and $\epsilon$ in the reconstruction, the average purity increased to $\bar{P}=0.995$ and the minimum purity increased to $0.982$.
-The information gained in the calibration clearly mitigated the reconstruction artifact.
-As a next step, we use this information to improve state preparation and measurement control.
-We calculated the new angular position of waveplates, considering the discovered retardance deviations, to implement the improved preparation and projection and repeated the experiment, obtaining $F_{\min}=0.991$, $\bar{F}=0.997$, $\bar{P}=0.995$, and $P_{\min}=0.983$.
-The results shown in Figure 2 indicate improved quality of state preparation and tomographic measurements.
+We minimize the apparent purity modulation from 0.0700(3) to 0.0144(2) to find the optimal parameters $\tilde{\delta}=4.53(6)$ deg and $\tilde{\epsilon}=-3.58(3)$ deg.
+The calibration step revealed the retardance deviation of the wave plates used for preparation and projection.
+To reconstruct the quantum states, we could directly use this knowledge to update our measurement operators.
+However, we take one extra step which allows us to prepare the states with greater precision and control the measurement operators precisely.
+We took these deviations into account in the verification step and calculated new angular positions of wave plates for tomographic projections and state preparation.
+Since Eq.
+10 does not hold anymore, we used numerical optimization to find the desired wave plate settings ($x,y$):
+$x,\,y=\mathrm{argmax}_{x,y}|\langle\psi|\hat{W}(x,\pi+\delta)\hat{W}(y,\pi/2+\epsilon)|0\rangle|^{2}.$ (11)
+Similarly, to find the settings to project a state onto state $|\pi\rangle$ we perform this maximization
+$x,\,y=\mathrm{argmax}_{x,y}|\langle\langle 0|\hat{W}(x,\pi/2+\epsilon)\hat{W}(y,\pi+\delta)|\pi\rangle|^{2}.$ (12)
+At this point, we have information about actual measurement operators, and we have also calibrated the state preparation.
+This calibration was done in the presence of a unitary operation between state preparation and measurement, which was introduced by QWPX.
+This operation was effectively unknown because we did not use any prior information about QWPX during the process.
+Now, we proceed to certify the correctness of such a calibration.
+The calibration should remain valid even if we change the central unitary operation introduced by QWPX.
+To test this, we removed the QWPX and then prepared an ensemble of 58 test states, depicted as empty circles in Figure 2(c,d).
+As a reference, we first assumed perfect waveplates and performed state tomography for each test state.
+We indeed observed purity modulation $\Delta P=0.0607(2)$ with purity minimum $P_{\min}=0.9383(2)$.
+This modulation indicates reconstruction artifacts.
+The minimum fidelity of the reconstructed state to the target states was $F_{\min}=0.96688(4)$.
+These fidelity values indicate imprecise preparation.
+In contrast, when we used our knowledge of $\delta$ and $\epsilon$ in the reconstruction and preparation, the minimum purity increased to $0.9830(3)$ and the minimum fidelity to $F_{\min}=0.9907(2)$.
+The information gained in the calibration clearly improved the reconstruction.
+The results in Figure 2 indicate improved quality of state preparation and tomographic measurements when we use the knowledge of $\delta$ and $\epsilon$ to improve both preparation and measurement.
 The small remaining artifacts are present due to the limited precision of the learned retardance deviations and also the repeatability of the waveplate rotation.
-Further, in Section S4 of the Supplementary material, we show the use of the presented method for classical rotating waveplate polarimetry [6].
-Conclusion - The mismatch between actual measurement operators and their theoretical counterparts assumed in reconstructing quantum states or devices introduces reconstruction artifacts.
+VII Path-encoded qubits on a photonic chip
+Figure 3: Application of the method for on-chip photonic path qubit.
+(a) Projection apparatus for a path qubit.
+Applied voltage $V_{1,2}$ on heater elements (orange) causes thermal-based phase shift in the waveguide interferometers.
+These shifts determine the measurement projector.
+The photons in the upper output path are detected with a single-photon detector (SPAD).
+(b) 1D-section of 6D function $\Delta P(c_{1},\dots,c_{6})$.
+The value of $\Delta P$ at the initial assumption of $c_{j}$ is $1.8\cdot 10^{-2}$ on the ensemble of probe states.
+Minimum $\Delta P=4.5\cdot 10^{-4}$ lies at the true value $c^{\prime}_{6}$.
+(c) Initial mismatch results in state-dependent purity artifact.
+We plot test states on the Bloch sphere as circles and color-code their reconstructed purity.
+(d) When we consider the true retardance, the artifact vanishes.
+As another example, we numerically illustrate the measurement-device agnostic calibration for tomography of photonic on-chip path qubits [30, 35, 36, 37].
+The qubit is encoded into the photon propagation in one or another waveguide (states $\ket{\uparrow}$ or $\ket{\downarrow}$), as depicted in Figure 3(a).
+The local heating of one waveguide controls the phase between the two paths.
+The heating is typically achieved by an electrical current flowing through a small heater element.
+A Mach-Zehnder interferometer serves as a variable coupler, and its interferometric phase controls the coupling ratio.
+Detection of photons in a single optical path implements a projective measurement of the path qubit.
+The circuit in Figure 3(a) implements a projection onto state |$\ket{\pi}$, whose colatitude and longitude on the Bloch sphere are controlled by optical phases.
+The relation between optical phases $\varphi_{1,2}$ and voltages $V_{1,2}$ applied on the heating elements is approximated with polynomials [35]
+$\varphi_{1}(V_{1})=c_{1}+c_{2}V_{1}+c_{3}V_{1}^{2},$ (13)
+$\varphi_{2}(V_{2})=c_{4}+c_{5}V_{2}+c_{6}V_{2}^{2}.$ (14)
+Here, the voltages $V_{1,2}$ are controlled perfectly, but parameters $c_{j}$ are known only approximately.
+The other imperfections are neglected.
+We have increased the number of probe states to 22 to compensate for the higher number of optimization parameters.
+The simulated experiment, depicted in Fig.
+3(c), shows the measurement operator mismatch artifact.
+We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j})$ where the systematic error is mitigated.
+Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c^{\prime}_{6}$.
+Taking the true values $c^{\prime}_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).
+VIII Conclusion
+The mismatch between actual measurement operators and their theoretical counterparts assumed in reconstructing quantum states or devices introduces reconstruction artifacts.
 We showed how to leverage these artifacts to reveal the actual measurement operators and mitigate systematic errors in the tomography and state preparation.
@@
 in the case of integrated circuits.
-One can also calibrate the state preparation with the reversed version of the method.
-Furthermore, the presented approach applies locally on individual parties of a larger multiparty system; i.e.
+One can also calibrate the state preparation using the same method.
+Furthermore, the presented approach applies locally to individual parties of a larger multiparty system; i.e.
 it is fully scalable.
 In summary, we developed the framework for accurate and scalable calibration of quantum measurement devices and also quantum state preparation, which is critical in any applications of quantum science and technology.
-The application of the reported measurement-device agnostic quantum measurement goes beyond the full quantum tomography.
-The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches such as Monte-Carlo sampling [36, 37], permutationally invariant tomography [38], compressed sensing [39, 40], etc.
+The application of the reported measurement-device agnostic quantum measurement goes beyond the full quantum tomography [13, 22].
+The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches such as Monte-Carlo sampling [38, 39, 40, 41], permutationally invariant tomography [42], compressed sensing [43, 44], etc.
 The method can also improve the accuracy of building and characterization of the photonic and quantum circuits in complex media [45, 46].
@@
 Acknowledgements.
-Acknowledgments - We acknowledge the support of the Czech Science Foundation under grant No.
+We acknowledge the support of the Czech Science Foundation under grant No.
 21-18545S.
@@
 Provazník for maintaining the cluster and providing support.
-Supplemental Material: Measurement-device agnostic quantum tomography
-I NOMAD setting in the numerical test
-When we numerically investigated the case of general additive errors, we set the NOMAD [1] parameters as listed in Tab.
-1.
-Please refer to the NOMAD 4 manual [2] for the meaning of the parameters.
-The optimized parameters were bounded to the interval (-29, 29) deg, i.e., roughly three standard deviations of the true parameter distribution.
-The seed was pseudo-randomly chosen at the time of optimization.
-The maximum size of the evaluation block was set as twice the problem dimension to allow optimal parallelization of the initial Latin hypercube search.
-We used the nomadlad [3] package to use NOMAD from Python.
-The necessary configuration was to set NOMAD to use multiple initial optimization points instead of just one by not specifying parameter X0 and setting the Latin hypercube search parameter (LH SEARCH) sufficiently high.
-parameter value
-BB OUTPUT TYPE OBJ
-DIMENSION 12
-BB MAX BLOCK SIZE 24
-MAX BB EVAL 100 000
-SEED random integer
-DIRECTION TYPE ORTHO 2N
-LH SEARCH 1100 1100
-LOWER BOUND (-0.5, …, -0.5)
-UPPER BOUND (0.5, …, 0.5)
-Table 1: NOMAD (v4.1.0) settings used for the global optimization. Some parameters are not listed because they are not critical for this application.
-II Artifact mitigation of multiplicative errors
-Here, we explore the sampling strategies in the case of multiplicative errors.
-Since the artifact is state-dependent, the probe sampling strategy is important for minimizing the reconstruction artifacts.
-As in the main text, all states are measured in projections onto eigenstates of Pauli operators.
-The projections are implemented by rotating the measured state using the following unitary operation $\hat{R}_{y}(\theta_{j})\hat{R}_{z}(\phi_{j})$ and subsequent projection of the resulting state onto state $|0\rangle$.
-The rotation angles $\theta_{j},\phi_{j}$ determine the selected projection and are listed in Tab 2.
-We test a few sampling schemes and plot the cost function near the ground truth ($\delta^{\prime}=0.02$, $\epsilon^{\prime}=-0.04$) for each.
-The cost function is the peak-to-valley value of reconstructed probe state purities, $\Delta P=P_{\max}-P_{\min}$.
-The probe sets and cost functions are depicted in Figure S1.
-As probe states set, we selected eigenstates of Pauli operators (panels a,b), eight states in (panels c, d), icosahedron (panels e, f), and then a uniform sampling of Bloch sphere angular coordinates ($\theta,\phi$) with 14 states (panels g, h) and 22 states (panels i, j) and finally NASA HEALPIX [4, 5] with 108 probe states (panels k, l).
-$j$ 1 2 3 4 5 6
-$\theta_{j}$ ($y$-axis) 0 $\pi$ $\pi/2$ $\pi/2$ $\pi/2$ $\pi/2$
-$\phi_{j}$ ($z$-axis) 0 0 $\pi$ 0 $\pi/2$ $23\pi/$ 2
-Table 2: Rotation angles in radians for Pauli tomography.
-The analyzed state is first rotated by $\phi_{j}$ around the $z$-axis and $\theta_{j}$ around the $y$-axis and subsequently projected to state $|0\rangle$.
-The sampling is sufficient when the cost function possesses a unique global minimum.
-We have run additional simulations, varied the ground truth, and observed the shape of the cost function.
-We tested each sampling scheme from Fig.
-S1.
-We sampled the cost function near the varying ground truth values of $\delta^{\prime}$ and $\epsilon^{\prime}$ and plotted its value as color maps, depicted in Fig.
-S1.
-The first suitable sampling scheme appears for eight probe states.
-Interestingly, for 14 probe states, there is a rift in the cost function landscape that causes ambiguity in the $\delta,\epsilon$ determination.
-The remaining sampling schemes exhibit a well-defined global minimum and therefore are suitable.
-The cost function landscapes in Fig.
-S1 are plotted for a single choice of $(\delta,\epsilon)$.
-To better characterize the sampling strategies, we repeated the same simulations for multiple ground truths.
-Figures S2-S6 show the simulated cost function for 8, 12, 17, 22, and 108 probe states.
-The ground truth is always plotted in the panel center.
-This additional characterization confirms the result – 8 probe states are enough and there is indeed anomalous behavior for 14 probe states.
-For 22 states, we sometimes see a local minimum, such as for ground truth parameters $\delta^{\prime}=0.15$, $\epsilon^{\prime}=0$.
-Due to this, the optimization has to be done carefully, ideally with several starting points, to find the global minimum of the cost function.
-In the case of 108 probe states, the cost function always had a single minimum.
-The most interesting outcome of this analysis is that the success of the sampling strategy is not determined by the sheer number of probe states but rather by their suitable placement on the Bloch sphere.
-Generally, adding more probe states makes the error parameters determination more robust because we increase the obtained information, however, at the cost of longer measurement time.
-The exact cost function shape, sampling strategies, and shapes of the purity modulation depend on the imperfection type, its parameterization, and how we switch the measurement operators in the tomography.
-Figure S1: Dependence of simulated artifact severity, represented as purity modulation amplitude depicted as 2D color maps in panels (a, c, …, k) on the selection of probe states set, depicted as points on projected Bloch sphere in panels (b, d, …, l).
-The figure shows the case for 6 (a,b), 8 (c,d), 12 (e,f), 14 (g,h), 22 (i,j), and 108 (k,l) probe states.
-The purity modulation amplitude is always plotted around the ground truth $\delta=0.02$, $\delta=-0.04$.
-Figure S2: The panels show color-coded purity modulation as a function of assumed error parameters ($\delta,\epsilon$) observed in the set of eight reconstructed probe states.
-The purity modulation is sampled near the true value ($\delta^{\prime},\epsilon^{\prime}$) for each simulated ground truth.
-This true value is always plotted in the center of the panel.
-All panels share the color bar.
-To visually emphasize the presence of a single minimum, we use logarithmic normalization of color coding.
-Figure S3: 12 probe states. The figure is arranged as Fig. S2.
-Figure S4: 14 probe states. The figure is arranged as Fig. S2.
-Figure S5: 22 probe states. The figure is arranged as Fig. S2.
-Figure S6: 108 probe states. The figure is arranged as Fig. S2.
-III Experimental setup details
-We used a continuous-wave 810 nm fiber-coupled laser diode of 1 mW power.
-The laser diode was thermally stabilized using Peltier elements.
-The laser beam was decoupled into free space using an 11-mm aspheric lens and horizontally polarized using a Glan-Taylor prism.
-The angular positions of quarter-wave and half-wave plates controlled the resulting polarization of the beam.
-The following operator describes the action of a waveplate:
-$\hat{W}_{\alpha,\gamma}=|\alpha\rangle\langle\alpha|+e^{-i\Gamma}|\alpha_{\bot}\rangle\langle\alpha_{\bot}|,$ (S1)
-where $\alpha$ and $\Gamma$ are its angular position and retardance, and $|\alpha\rangle=\cos{\alpha}|0\rangle+\sin{\alpha}|1\rangle$ is a linearly polarized state, and $\alpha_{\bot}=\alpha+\pi/2$.
-To prepare a state located at colatitude $\theta$ and longitude $\phi$ on the Bloch sphere, we rotate the half- and quarter-wave plate to angles $x$ and $y$ given by the relations
-$y=-\frac{1}{2}\arcsin\left(\sin\theta\sin\phi\right),$ (S2)
-$x=\frac{1}{4}\arctan\left(\tan\theta\cos\phi\right)+\frac{1}{2}y+c(\theta),$ (S3)
-where $c(\theta)=\pi/4$ if $\theta>\pi/2$ and $c(\theta)=0$ otherwise.
-This relation holds for wave plates that have perfect retardance.
-The deviation from the ideal retardance will introduce errors in the preparation.
-We assumed the perfect retardance in the first step of the experiment as our method does not require exact knowledge of probe states.
-Then we inserted a rotated quarter-wave plate that rotated the prepared state to demonstrate the robustness of our method to unitary operations in between the preparation and projection stages.
-The states were analyzed using half-waveplate and quarter-wave and subsequent projection onto horizontal and vertical polarization using a calcite beam-displacer.
-The waveplate settings for the Pauli tomography are listed in Tab.
-3.
-All waveplates were mounted into a precision motorized rotation stage that provided bidirectional repeatability of ±$\pm 30$ mdeg.
-The optical signals were measured by reading photo-currents from a pair of reverse-biased photodiodes.
-We normalized the readings using the sum of two signals to eliminate laser power fluctuation.
-The same results are obtained for a single-photon signal using two single-photon detectors.
-The calibration step revealed the retardance deviation of the preparation and projection wave plates.
-To reconstruct the quantum states, we could directly use this knowledge to update our measurement operators.
-However, we take one extra step which allows us to prepare the states with greater precision and control the measurement operators precisely.
-We took these deviations into account in the verification step and calculated new angular positions of wave plates for tomographic projections and state preparation.
-Since Eq.
-S3 does not hold anymore, we used numerical optimization to find the desired wave plate settings.
-To prepare a state $|\psi\rangle$, we find the wave plate angles $x,y$ using the following maximization:
-$x,y=\mathrm{argmax}_{x,y}|\langle\psi|\hat{W}(x,\pi+\delta)\hat{W}(y,\pi/2+\epsilon)|0\rangle|^{2}.$ (S4)
-Similarly, to find the settings to project a state onto state $|\pi\rangle$ we perform this maximization
-$x,y=\mathrm{argmax}_{x,y}|\langle\langle 0|\hat{W}(x,\pi/2+\epsilon)\hat{W}(y,\pi+\delta)|\pi\rangle|^{2}.$ (S5)
-state $|H\rangle$ $|V\rangle$ $\frac{|H\rangle+|V\rangle}{\sqrt{2}}$ $\frac{|H\rangle-|V\rangle}{\sqrt{2}}$ $\frac{|H\rangle+i|V\rangle}{\sqrt{2}}$ $\frac{|H\rangle-i|V\rangle}{\sqrt{2}}$
-$j$ 1 2 3 4 5 6
-$x_{j}$ (HWP) 0 $\pi/4$ $\pi/8$ $-\pi/8$ $-\pi/8$ $\pi/8$
-$y_{j}$ (QWP) 0 0 0 0 $-\pi/4$ $\pi/4$
-$x^{\prime}_{j}$ (HWP) 0 $\pi/4$ $\pi/8$ $-\pi/8$ $\pi/8$ $-\pi/8$
-$y^{\prime}_{j}$ (QWP) 0 0 0 0 $\pi/4$ $-\pi/4$
-Table 3: Waveplate angular position for Pauli tomography.
-The waveplates are placed as follows in the direction of the light propagation: quarter-wave plate and half-wave plate in the preparation, and half-wave plate and quarter-wave plate in the measurement stage.
-Primed symbols correspond to the state preparation, used in the reversed version of the method.
-IV Rotating waveplate polarimetry
-Consider the classical rotating quarter-wave plate polarimeter [6], depicted in Fig.
-S7(a), with the quarter-wave plate having the true retardance deviated by $\delta^{\prime}=5\deg$ from its nominal value.
-Moreover, consider pure probe states.
-The waveplate rotates continuously and effectively modulates the optical signal detected after the polarizer.
-Instead of likelihood maximization, we now use classical Fourier analysis to reconstruct the polarization state.
-We chose the Fourier analysis to show that our method is not limited to maximal likelihood reconstruction.
-We used eight probe states to detect the reconstruction artifact, and instead of purity modulation, we equivalently measured the severity with a modulation amplitude of the degree of polarization.
-First, we treat the data with the ideal retardance ($\delta=0$) assumption.
-We see that the states from the ensemble suffer from artifacts.
-Here, we look at the intensity-normalized Stokes vector length $D=\frac{\sqrt{\sum\limits_{i=x,y,z}S_{i}}}{S_{0}}$, which is proportional to purity $P=\frac{1}{2}\left(1+D^{2}\right)$.
-Near state |$\ket{H}$, at the north pole of the Bloch sphere, the purity even exceeds 1 due to the reconstruction method.
-The modulation of $D$ is depicted in Fig.
-S7(c).
-This behavior is clearly a reconstruction artifact.
-Exceeding the unit purity is unsurprising because Fourier analysis does not guarantee physically sound results.
-We now vary $\delta$ assumed in reconstruction and observe the peak-to-peak value $\Delta D$ in the probe set.
-The dependence is depicted in Fig.
-S7(b).
-We see that the minimum lies at the true $\delta^{\prime}$.
-In panel (d), we show that the artifact disappears when the retardance deviation is taken into account.
-Figure S7: Mitigation of artifact in rotating-waveplate polarimetry.
-(a) Measurement apparatus.
-A quarter-wave plate (purple) with retardance deviation $\delta^{\prime}=5$ deg is rotating in front of the linear polarizer.
-The photodiode signal is recorded and analyzed.
-(b) The minimal $\Delta D$ lies at the true retardance deviation $\delta^{\prime}$.
-(c) Initial mismatch results in state-dependent purity artifact.
-We plot the reconstructed probe states on the projected Bloch sphere and color-code their reconstructed stokes length $D=\sqrt{S^{2}_{x}+S^{2}_{y}+S^{2}_{z}}/S_{0}$.
-(d) When we considered the true retardance, the artifact vanished.
-Black circles depict the ideal testing states.
-V Simulated experiment with path-encoded qubits on a photonic chip
-As another example, we illustrate the measurement-device agnostic calibration for tomography of photonic on-chip path qubits [7, 8, 9, 10].
-The qubit is encoded into the photon propagation in one or another waveguide (states $\ket{\uparrow}$ or $\ket{\downarrow}$), as depicted in Figure S8(a).
-The local heating of one waveguide controls the phase between the two paths.
-The heating is typically achieved by an electrical current flowing through a small heater element.
-A Mach-Zehnder interferometer serves as a variable coupler, and its interferometric phase controls the coupling ratio.
-Detection of photons in a single optical path implements a projective measurement of the path qubit.
-The circuit in Figure S8(a) implements a projection onto state |$\ket{\pi}$, whose colatitude and longitude on the Bloch sphere are controlled by optical phases.
-The relation between optical phases $\varphi_{1,2}$ and voltages $V_{1,2}$ applied on the heating elements is approximated with polynomials [8]
-$\varphi_{1}(V_{1})=c_{1}+c_{2}V_{1}+c_{3}V_{1}^{2},$ (S6)
-$\varphi_{2}(V_{2})=c_{4}+c_{5}V_{2}+c_{6}V_{2}^{2}.$ (S7)
-Figure S8: Application of the method for on-chip photonic path qubit.
-(a) Projection apparatus for a path qubit.
-Applied voltage $V_{1,2}$ on heater elements (orange) causes thermal-based phase shift in the waveguide interferometers.
-These shifts determine the measurement projector.
-The photons in the upper output path are detected with a single-photon detector (SPAD).
-(b) 1D-section of 6D function $P_{\min}(c_{1},\dots,c_{6})$.
-The value of $P_{\min}$ at the initial assumption of $c_{j}$ is 0.955.
-Maximum lies at the true value $c^{\prime}_{6}$.
-(c) Initial mismatch results in state-dependent purity artifact.
-We plot probe states on the Bloch sphere and color-code their reconstructed purity.
-(d) When we consider the true retardance, the artifact vanishes.
-Parameters of the simulation are: $c^{\prime}_{1}=-1.887$, $c^{\prime}_{2}=0.105$ V-1, $c^{\prime}_{3}=0.05$ V-2, $c^{\prime}_{4}=-1.805$, $c^{\prime}_{5}=0.115$ V-1, and $c^{\prime}_{6}=0.05$ V-2.
-Initially assumed parameters were: $c_{1}=-1.750$, $c_{2}=0.100$ V-1, $c_{3}=0.050$ V-2, $c_{4}=-1.800$, $c_{5}=0.1105$ V-1, and $c_{6}=0.050$ V-2
-Here, the voltages $V_{1,2}$ are controlled perfectly, but parameters $c_{j}$ are known only approximately.
-The other imperfections are neglected.
-We have increased the number of probe states to 22 to compensate for the higher number of optimization parameters.
-The simulated experiment, depicted in Fig.
-S8(c), shows the measurement operator mismatch artifact.
-We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j})$ where the systematic error is mitigated.
-Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c^{\prime}_{6}$ (indicated by the dashed line).
-Taking the true values $c^{\prime}_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).

### v5

@@
 Characterization of quantum states and devices is paramount to quantum science and technology.
-The characterization consists of individual measurements, which are required to be precisely known.
+The characterization consists of individual measurements, which must be precisely known.
 A mismatch between actual and assumed constituent measurements limits the accuracy of this characterization.
-Here, we show that such a mismatch introduces reconstruction artifacts in quantum state tomography.
-We use these artifacts to detect and quantify the mismatch and gain information about the actual measurement operators.
-It consequently allows the mitigation of systematic errors in quantum measurement and state preparation.
+We show that such a mismatch introduces reconstruction artifacts in quantum state tomography.
+We use these artifacts to detect and quantify the mismatch, gaining information about the actual measurement operators.
+It consequently allows the mitigation of systematic errors in both quantum measurement and state preparation, improving the precision of state control and characterization.
+The practical utility of our approach is experimentally demonstrated.
 I Introduction
-The measurements play a central role in quantum science and technology [1].
-However, the implemented measurements often differ from the ideal ones, leading to systematic errors.
-One can compensate for these systematic errors with the knowledge of the actual measurements.
-We investigate this problem from the point of view of quantum tomography, one of the most valuable tools in quantum science [2, 3, 4, 5, 6].
+Measurement plays a central role in quantum science and technology [1].
+However, the implemented measurements often systematically differ from the desired ones.
+Such a discrepancy can be understood as a systematic error.
+The main focus of this work is on detecting and mitigating these systematic errors.
+Mitigating systematic error also means gaining information about the measurement operators, which can be used to determine a proper compensation that matches the actual measurements with the desired ones.
+We investigate this problem from the perspective of quantum tomography, one of the most valuable tools in quantum science [2, 3, 4, 5, 6].
 It allows the complete characterization of prepared states and implemented quantum circuits, assessing and certifying their quality.
@@
 One can reconstruct the state from the collection of measurement outcomes, the tomogram.
-Imperfect realization of these measurements causes a mismatch between the assumed and the realized measurement operators.
+The imperfect realization of these measurements causes a mismatch between the assumed and the realized measurement operators.
 Such a mismatch leads to the wrong interpretation of the acquired tomogram and manifests as a reconstruction artifact.
-The true measurement operators can be fully characterized with measurement tomography [7, 8, 9, 10], which relies on perfect knowledge of input probe quantum states.
-Such a requirement is difficult to meet in the experiments.
-A self-consistent tomography framework has been introduced to loosen this condition [11].
-The framework suggests the existence of a self-calibrating state class, which could reveal, when measured, the initially unknown information about the states and the measurement device.
-Another approach demonstrated for polarization states of light infers a single unknown parameter of the measurement device using maximum likelihood estimation [12].
-There are also schemes that use simultaneous likelihood maximization over both states and devices [13] or states and measurements [14].
-Finally, assuming the sparsity of the parameter vector and the low rank of the measured density matrix, the dependence on the measurement device can be relaxed [15].
-In this work, we show that an ensemble of single-qubit states of equal purity distributed quasi-uniformly on the Bloch sphere forms a class of self-calibrating states.
-This class is convenient in experiments where states are naturally prepared with quasi-equal purity, e.g., in photonics and trapped ions.
-Advantageously, the precise control, characterization, or even sparsity of such self-calibrating probe states is not required.
-We show a novel approach to simultaneously calibrating quantum measurements and state preparation.
+Systematic errors in quantum experiments often originate from inaccurate assumptions about the properties or performance of preparation and measurement devices.
+Such errors may arise from imperfect knowledge of device parameters, deviations of operating conditions from their nominal values, or limitations of prior calibration procedures.
+For example, the retardance of birefringent wave plates in photonic devices is typically assumed to be exactly $\pi$ (half-wave) or $\pi/2$ (quarter-wave), yet manufacturing tolerances and wavelength mismatches can shift the actual retardance by a few degrees, leading to systematic misrotation of measurement bases.
+Similarly, in trapped-ion experiments, miscalibrations of the laser intensity or pulse length alter the Rabi frequency, producing consistent over- or under-rotations of qubits.
+These and related discrepancies between assumed and actual device behavior give rise to systematic errors (bias) that persist across repeated measurements.
+The true measurement operators can be fully characterized with measurement tomography [7, 8, 9, 10].
+While in standard quantum state tomography, we use perfectly known measurements to reconstruct a completely unknown state, in measurement tomography, we use perfectly known states and measure them using an unknown detector, eventually reconstructing the operator that described the measurement.
+Such a requirement is challenging to meet in experiments, as state preparation can be affected by imperfections and miscalibrations, as well as detection.
+Albeit the existence of a scheme for simultaneous reconstruction of a quantum channel with unknown states [11], the challenge remains.
+The scheme still requires perfectly known measurements and access to input states before they are affected by the channel.
+One could model the systematic error in the measurement as a fixed unknown channel placed between the state preparation and the detector.
+However, this approach would be very limiting and would not cover cases where the error varies for each measurement operator used in state tomography.
+A self-consistent tomography framework has been introduced to loosen the condition of having perfectly known measurements [12].
+The framework suggests the existence of a self-calibrating state class, which, when measured, could reveal initially unknown information about the states and the measurement device.
+In the self-calibration process, one transfers the prior information about input states into information about the measurement device.
+In the original work, the authors provide an example of self-calibration using an on-off detector.
+Although the work provides a conceptual sketch of a generic self-calibration procedure, it was not applied to the system of qubits, i.e., the family of self-calibrating qubit states was not identified.
+The problem of quantum state tomography with imprecisely known measurement devices has been addressed [13] by utilizing perfect prior knowledge of a few (six) states used to calibrate the detection.
+Let us note that with such ideal input states, one could theoretically perform quantum measurement tomography [7], calibrate the detector, and then continue with state tomography.
+Therefore, such an approach does not address the issue of calibrating the measurement device with imperfectly known probe states.
+A work [14] addressed this challenge by identifying low-rank states as a family of self-calibrating states.
+The work provides a general and rigorous discussion of self-calibration.
+In its simplest form, it optimizes the reconstructed density matrices and the parameters describing systematic errors in the measurement to achieve the best fit of the model to the data.
+The maximum rank of the probe states must be less than an a priori known number.
+The error model describes the measurement outcome as a linear combination of perfect outcomes and outcomes corresponding to a changed measurement operator in a Pauli basis.
+The coefficients in the linear combination form an error vector.
+Another requirement of the method [14] is that the error vector has a known and high degree of sparsity.
+Therefore, the method is effective in cases where a unitary rotation about one of the three axes is dominant.
+So far, we have discussed theoretical works.
+An experimental demonstration of self-calibration [15] approaches this problem by introducing unknown parameters of the measurement device as another optimization parameter in the likelihood maximization.
+Then, with a single state tomography, one can recover the unknown parameter, retardance in this case.
+Since the measurement device parameter is estimated using a single state used to probe it, it is unclear whether there is some ambiguity and how to treat conflicting results about the parameters of a single measurement device calibrated by two different probe states.
+The work [15] also addresses the case of two qubits, by adding two unknown parameters into two-qubit tomography.
+Such an approach, however, would lead to an exponential increase in resources when scaling to more qubits.
+Another limitation is that the procedure studies only the case of multiplicative errors and does not discuss a more general class of errors.
+In this work, we show that an ensemble of single-qubit states of quasi-equal purity distributed quasi-uniformly on the Bloch sphere forms a class of self-calibrating states.
+This class is particularly convenient in experiments where states are naturally prepared with quasi-equal purity, such as in photonics and trapped ions.
+Advantageously, the precise control or characterization of such self-calibrating probe states is not required.
+We present a novel approach to calibrating quantum measurements and state preparation within a single framework, eliminating the need for physical rearrangements of the experiment.
 To this end, we exploit artifacts in quantum state reconstruction performed on the self-calibrating probe states.
-In the multi-qubit setting, we can use the method in a fully scalable way to mitigate systematic errors on each local qubit measurement, effectively mitigating errors in the whole system.
-The method also applies to quantum device tomography thanks to the state-process duality [16, 17, 1], which allows the quantum processes to be treated as states.
-Our method is not limited to the determination of an initially unknown parameter shared among all measurement operators in the tomographic scheme; it allows the determination of parameters related to each measurement operator individually.
+We also assess how much noise our method can tolerate and still provide an advantage over the calibration-less approach.
+In the multi-qubit setting, we can utilize the method in a fully scalable manner to mitigate systematic errors in each local qubit measurement, thereby effectively reducing errors throughout the entire system.
+The method also applies to quantum device tomography thanks to the state-process duality [16, 17, 1], which allows treating quantum processes as states.
+Our method is not limited to the determination of an initially unknown parameter shared among all measurement operators in the tomographic scheme; it also allows the determination of parameters related to each measurement operator individually.
 The presented scheme provides the characterization of quantum states independently of measurement device imperfections.
-General concepts are introduced using numerical simulations, and we support our findings with a photonic experiment.
+General concepts are introduced using numerical simulations, and we support our findings with a photonic experiment, surpassing the fidelity reported for measurement self-calibration in previous works [15].
+Moreover, the presented method does not require modification of the used reconstruction algorithm - it can be employed as a black-box function in our procedure.
 Figure 1: Purity modulation due to the measurement operator mismatch.
-(a-c) States are represented as points on the Bloch sphere plotted using Hammer projection, and the purity of their reconstruction is color-coded.
+(a-c) States are represented as points on the Bloch sphere plotted using Hammer projection, and the purity of their reconstructions is color-coded.
 Although all probe states were prepared with the same purity, the reconstruction using quantum tomography with mismatched measurement operators yields purity modulation.
@@
 II Measurement operator mismatch artifact
-In quantum state tomography, many copies of an initially unknown state, described by density operator $\hat{\rho}$, are subject to a series of measurements described by measurement operators $\hat{\pi}_{j}$.
-These measurement operators should be tomographically complete, and their choice influences the performance of the tomography [18, 19, 20, 21].
+As a first step, let us investigate how the discrepancies between assumed and actual measurement operators manifest in reconstruction artifacts.
+In quantum state tomography, many copies of an initially unknown state, described by density operator $\hat{\rho}$, are subject to a series of measurements described by measurement operators $\hat{\pi}_{j}$ drawn from a set of positive-operator-valued measure (POVM), which are positive and add up to the identity operator [1].
+These measurement operators should be tomographically complete, i.e., the density operator must be decomposable into their linear combination [1], and their choice influences the performance of the tomography [18, 19, 20, 21].
 The resulting collection of measurement outcomes forms the tomogram.
 The tomogram and the measurement operators are then input into an algorithm, which outputs the density operator.
+Let us note that the measurement operators are a necessary input to the reconstruction algorithm, as they provide crucial context for interpreting the tomogram [22].
 Here, we use the maximum-likelihood reconstruction algorithm [11, 22].
-If the actual measurement operators $\hat{\pi}^{\prime}_{j}$ differ from those put into the reconstruction method, artifacts appear.
-Let us demonstrate the behavior in the simplest scenario of pure probe single-qubit states and projective measurements onto the eigenstates of Pauli operators, called Pauli tomography.
-The measurement operator is realized by two rotation angles, corresponding to rotation about the $y$- and $z$-axis and subsequent projection onto computational basis state $|0\rangle\langle 0|$.
+If the actual measurement operators $\hat{\pi}_{j}$ differ from those used in the reconstruction method, $\hat{\pi}^{\prime}_{j}$, that is $\hat{\pi}_{j}\neq\hat{\pi}^{\prime}_{j}$, artifacts appear.
+To demonstrate the behavior in the core principle of our method, we use pure probe single-qubit states and projective measurements onto the eigenstates of Pauli operators, called Pauli tomography.
+We must state that this choice is arbitrary and the presented framework is independent of a particular POVM choice.
+The projective measurement operators are realized by two rotation angles, corresponding to rotation about the $y$- and $z$-axis and subsequent projection onto computational basis state $|0\rangle\langle 0|$.
 Throughout the article, we use the computational basis $\{|0\rangle,|1\rangle\}$.
@@
 $p_{j}=\mathrm{Tr}[\hat{\rho}\hat{\pi}_{j}].$ (2)
-The angles $\theta_{j}$ and $\phi_{j}$ for Pauli tomography are provided in Table S2 of the Supplemental Material. Assume that due to experimental imperfections, the actual angles are
-$\theta_{j}^{\prime}=\theta_{j}+\delta_{j},$ (3)
-$\phi_{j}^{\prime}=\phi_{j}+\epsilon_{j},$ (4)
+The angles $\theta_{j}$ and $\phi_{j}$ for Pauli tomography are provided in Table S2 of the Supplemental Material. Assume that due to systematic experimental imperfections, the actual angles are
+$\tilde{\theta}_{j}=\theta_{j}+\delta_{j},$ (3)
+$\tilde{\phi}_{j}=\phi_{j}+\epsilon_{j},$ (4)
 i.e., we introduce systematic additive errors for each measurement operator.
-The error vector $\vec{\delta}=\{\delta_{j},\epsilon_{j}\}$ parameterizes the measurement operators $\hat{\pi}^{\prime}_{j}=\hat{\pi}_{j}(\vec{\delta})$.
+We further assume that the parameters $\delta_{j}$ and $\epsilon_{j}$ remain effectively static throughout the calibration and evaluation, i.e., they do not exhibit stochastic fluctuations or temporal drift, or if they do, the rate of change is slow enough to be negligible during the calibration and evaluation period.
+We keep this assumption in the rest of the article.
+The error vector $\vec{\delta}=\{\delta_{j},\epsilon_{j}\}$ parameterizes the measurement operators.
 We numerically simulate tomograms using these true measurement operators for many pure probe states covering the Bloch sphere quasi-uniformly.
-Then, we reconstruct them assuming the original unperturbed measurement operators $\{\hat{\pi}_{j}\}$.
+Then, we reconstruct them assuming the original unperturbed measurement operators $\{\hat{\pi}^{\prime}_{j}\}$, i.e., with choice $\vec{\delta}^{\prime}=\vec{0}$.
 This mismatch causes the state-dependent purity modulation, shown in Figure 1(a-c).
@@
 where $P=\mathrm{Tr}\left(\hat{\rho}^{2}\right)$ is the purity of the reconstructed state.
-To show how purity modulation scales with the error vector magnitude, we scale it with a non-negative real factor $\lambda$, i.e.
-$\vec{\delta}\rightarrow\lambda\vec{\delta}$, and observe the $\Delta P$, as shown in Figure 1(d).
+To show how purity modulation scales with the error vector magnitude, we scale it with a non-negative real factor $\lambda$, i.e., $\vec{\delta}\rightarrow\lambda\vec{\delta}$, and observe the $\Delta P$, as shown in Figure 1(d).
 We see that $\Delta P$ monotonously increases with $\lambda$.
@@
 Moreover, the shape of the purity landscape is determined by the error vector $\vec{\delta}$.
-We will leverage this dependence of apparent purity as a function of vector $\delta$ to find the true measurement operators.
-The method is not limited to the particular reconstruction method, as we illustrate in an example of rotating quarter-wave plate polarimetry with classical Fourier retrieval of Stokes parameters [23].
-The example is provided in Supplementary Material S4.
-In that case, the disagreement between actual and assumed measurement operators causes inconsistency in the tomogram, which results in apparent state-dependent modulation of the degree of polarization, where the degree of polarization can even exceed one.
+This effect is not limited to the particular reconstruction method.
+In Supplementary Material S6, we provide an example of rotating quarter-wave plate polarimetry where we use classical Fourier retrieval of Stokes parameters [23] as a reconstruction method.
+The disagreement between actual and assumed measurement operators causes inconsistency in the tomogram, which results in apparent state-dependent modulation of the degree of polarization, where the degree of polarization can even exceed one.
+Apart from the purity modulation itself, other figures of merit, like fidelity, also scale with the error vector.
+For example, fidelity generally decreases with increasing $\lambda$.
+We, however, plot only the purity modulation because we intend to use it as the cost function for optimization.
+Since $\Delta P$ quantifies the severity of operator mismatch, we now leverage it to identify the true measurement operators and mitigate reconstruction artifacts.
 III Artifact mitigation
 We minimize the artifact severity, i.e., purity modulation (5), by varying the measurement operators used to reconstruct the probe states.
-After recording all tomograms of the probe states we used, we optimize our assumption of $\vec{\delta}$ which specifies the measurement operators used in the reconstruction algorithm.
-Once the assumed parameters match the actual parameters, the artifact is minimized.
-Since the perturbed measurement operators no longer add the identity operator, we must properly renormalize them in the reconstruction [22].
+After recording all tomograms of the probe states we used, we optimize our assumption $\vec{\delta}^{\prime}$, which specifies the measurement operators used in the reconstruction algorithm.
+We use a prime symbol to explicitly distinguish the assumed error vector $\vec{\delta}^{\prime}$ from the initially unknown actual error vector $\vec{\delta}$.
+Once the assumed parameters match the actual parameters, that is $\vec{\delta}^{\prime}\rightarrow\vec{\delta}$, the artifact is minimized.
+Since the perturbed measurement operators no longer add the identity operator, it is crucial to properly renormalize them in the reconstruction [22].
 Due to the state dependency, the Bloch sphere must be sampled densely enough to detect the mismatch.
@@
 It distinguishes our approach from the schemes that rely on perfect knowledge of the probe states, like quantum measurement tomography [8].
-Generally, the optimization landscape has multiple local minima; therefore, global optimization must be performed.
+Generally, the optimization landscape contains multiple local minima, and therefore global optimization must be performed.
+Optimizers that start from a single point and use gradient descent may converge to a local minimum, resulting in an incorrect error vector.
+Note that this issue is not related to whether or not the optimizer uses gradients but rather to whether the optimizer aims to find the global minimum.
 Here, we use the NOMAD implementation of the MADS algorithm [24].
-The resulting optimum $\vec{\delta}$ is not unique because any set of measurement operators $\{\hat{U}\hat{\pi}_{j}\hat{U}^{\dagger}\}$, where $\hat{U}$ is a fixed unitary operation, also minimizes the artifact.
+The resulting optimum $\vec{\delta}^{\prime}$ is not unique because any set of measurement operators $\{\hat{U}\hat{\pi}_{j}\hat{U}^{\dagger}\}$, where $\hat{U}$ is a fixed unitary operation, also minimizes the artifact.
 These unitary-equivalent solutions could be interpreted as unitary transformations of probe states.
-Albeit this ambiguity, such a calibration is already applicable in schemes invariant to local unitaries such as random measurement [25] or entanglement certification [26, 27, 28], where only relative orientations of $\hat{\pi}_{j}$ operators are important.
+Albeit this ambiguity, such a calibration is already applicable in schemes invariant to local unitaries, such as random measurement [25] or entanglement certification [26, 27, 28], where only relative orientations of $\hat{\pi}_{j}$ operators are important.
 Moreover, the unitary operation can be easily determined by additional measurement of two known and non-orthogonal probe states.
 Since the artifact is state-dependent, the probe sampling strategy is important for minimization.
-The sampling with too few probe states could lead to false error parameters $\vec{\delta}_{f}$.
+The sampling with too few probe states could lead to incorrect error parameters $\vec{\delta}^{\prime}_{f}$.
 We found out that 30 quasi-uniformly distributed probe states provide a reliable determination of error parameters in all tested cases.
@@
 Then, we produce tomograms corresponding to pure probe states $\rho_{k}=|\psi_{k}\rangle\langle\psi_{k}|$.
-For simplicity, we exclude statistical noise from our simulations.
-Parameters $\vec{\delta}^{\prime}$ assumed in the reconstruction are optimized to minimize the purity modulation $\Delta P$ in the set of reconstructed probe states $\{\hat{\rho}_{k}\}$ for $k=1,\dots,30$, concluding the calibration step.
-To verify the correctness of the optimization result, we use the ground truth $\vec{\delta}$ again to generate tomograms of pure test states $\hat{\xi}=|\xi_{k}\rangle\langle\xi_{k}|$, which quasi-uniformly covers the Bloch sphere in 108 points.
-We reconstruct these states using the optimized parameters $\vec{\delta}$ obtained in the calibration step and calculate the purity of the reconstructions.
+For simplicity, we exclude statistical noise from our simulations for now.
+Parameters $\vec{\delta}^{\prime}$ assumed in the reconstruction are optimized to minimize the purity modulation $\Delta P$ in the set of reconstructed probe states $\{\hat{\rho}_{j}\}$ for $j=1,\dots,30$, concluding the calibration step.
+To verify the correctness of the optimization result, we use the ground truth $\vec{\delta}$ again to generate tomograms of pure test states $\hat{\xi}=|\xi_{k}\rangle\langle\xi_{k}|$, which quasi-uniformly cover the Bloch sphere in 108 points.
+We reconstruct these states using the optimized parameters $\vec{\delta}^{\prime}$ obtained in the calibration step and calculate the purity of the reconstructions.
 As a further test, we assume knowledge of two non-orthogonal probe states and use their reconstructions to determine corrective unitary operation $\hat{W}$.
@@
 We characterize the test states ensemble with the purity modulation and their lowest fidelity $\min\limits_{k}F_{k}$ to provide the worst-case estimate.
-We repeated the whole numerical simulation 100 times, randomly choosing new ground truth for the error parameters $\delta$ in each run.
-The details about optimizer settings are provided in Section S1 of the Supplementary Material.
-When we compare the optimized results to the reference, i.e.
-the states reconstructed with the assumption of null error vector, the average purity modulation decreases from $\Delta P=(8\begin{subarray}{c}+3\\ -3\end{subarray}))\cdot 10^{-2}$ to $\Delta P=(1.0\begin{subarray}{c}+0.8\\ -1.0\end{subarray})\cdot 10^{-3}$.
-The uncertainty interval spans from 0.158 to 0.842-quantile in correspondence to one standard deviation and is used due to the skewed distribution of the results.
-Mainly, the infidelity decreased from $(3\begin{subarray}{c}+1\\ -2\end{subarray})\cdot 10^{-2}$ to $(3\begin{subarray}{c}+5\\ -3\end{subarray})\cdot 10^{-4}$, which represents two orders of magnitude improvement in accuracy.
+We repeated the entire numerical simulation 100 times, randomly drawing new ground truth for the error parameters $\delta$ in each run.
+The purpose of this randomization is to test the method performance on various possible combinations of deviations to assess its reliability.
+The details about optimizer settings are provided in Supplementary Material S1.
+When we compare the optimized results to the reference, i.e., the states reconstructed with the assumption of a null error vector, the average purity modulation decreases from $\Delta P=(8\begin{subarray}{c}+3\\ -3\end{subarray}))\times 10^{-2}$ to $\Delta P=(1.0\begin{subarray}{c}+0.8\\ -1.0\end{subarray})\times 10^{-3}$.
+The uncertainty interval spans from 0.158 to 0.842-quantile, corresponding to two standard deviations, and is used due to the skewed distribution of the results.
+Mainly, the infidelity decreased from $(3\begin{subarray}{c}+1\\ -2\end{subarray})\times 10^{-2}$ to $(3\begin{subarray}{c}+5\\ -3\end{subarray})\times 10^{-4}$, which represents a two orders of magnitude improvement in accuracy.
 This improvement is crucial for reaching error levels low enough for quantum error correction schemes.
-The lower $\Delta P$ of the probe state ensemble was achieved the better results in the test ensemble we achieved.
-These results illustrate the feasibility of such optimization and its ability to reliably determine the true measurement operators.
+The lower the $\Delta P$ of the probe state ensemble, the better the results we achieved in the test ensemble.
+A detailed scatter plot and marginal histogram of these numerical results can be found in Supplementary Material S2.
+The results illustrate the feasibility of the proposed calibration approach and its ability to reliably determine the true measurement operators.
 It is important to stress that the pure probe states and projective (pure) measurements were chosen here for the simplicity of introducing the measurement-device agnostic framework.
-However, the method is fully applicable in real cases of generic positive-operator-valued measurements and nonunity purity, as indicated in the following experimental demonstration.
-In the multi-qubit scenario, the tomography is usually done using local measurements.
-In that case, we use the presented method for each local measurement device.
-The artifact severity scales worst as $P_{\min,n}=\prod\limits_{i=1}^{n}P_{\min,i}$ for $n$-qubit state tomography, where $P_{\min,i}$ is the corresponding minimal purity observed at the single qubit.
-If the input state |$\ket{\psi_{i}}$ at the $i$-th qubit suffers from the artifact the most, then product state $\bigotimes\limits_{i}\ket{\psi_{i}}$ is the $n$-qubit state with the greatest manifestation of the artifact.
-In such a case, the single-qubit improvements multiply and lead to significantly better overall results in multi-qubit scenarios.
-IV State preparation
-The state preparation typically suffers from imperfections too.
-Here, we use the symmetry in the Born rule to formally exchange state preparation and measurement to utilize the previously introduced method to correct state preparation.
-Usually, the qubit is initialized in state $|0\rangle$, then turned into state $|\psi_{j}\rangle=\hat{U}_{j}|0\rangle$ by unitary evolution $\hat{U}_{j}$.
-The tomographic projection is also typically realized by some unitary evolution $\hat{V}_{k}$ and subsequent projection, i.e., $|\pi_{k}\rangle=\hat{V}_{k}^{\dagger}|0\rangle$.
-The state tomogram of the probe state $|\psi_{j}\rangle$ consists of measured probabilities $|\langle\hat{\pi}_{k}|\psi_{j}\rangle|^{2}=|\langle 0|\hat{V}_{k}\hat{U}_{j}|0\rangle|^{2}$.
-The discussed procedure obtains information related to projections, i.e.
-about $V_{k}$.
-Physical reversal of the process consists of preparing states $\hat{V}_{k}|0\rangle$ and projecting them onto states $\hat{U}_{j}^{\dagger}|0\rangle$, obtaining tomogram elements $p_{kj}\propto|\langle 0|\hat{U}_{j}\hat{V}_{k}|0\rangle|^{2}$.
-We can analyze the tomograms to improve our knowledge of the measurement operator, i.e., $\hat{U}_{j}^{\dagger}|0\rangle\langle 0|\hat{U}_{j}$, related to state preparation in the non-reversed situation.
-The physical reversal can be realized in the experiment by performing projections onto the original probe states in the tomographic device and preparing states corresponding to the original tomographical projectors instead of the original probes.
-Either way, we gain information about $\hat{U}_{j}$, increasing the accuracy of state preparation.
-V Special case: multiplicative error
+However, the method is fully applicable in real cases of generic POVM measurements and nonunity purity, as indicated in Section VI using the experimental demonstration.
+To further assess the method robustness, we investigated its performance under two practical noise sources: statistical shot noise and environmental decoherence.
+In the presence of finite measurement statistics, the method begins to outperform the calibration-less case with $10^{3}$ shots per measurement setting, and closely approaches the ideal (infinite-shot) performance at around $10^{5}$ shots.
+These results can be expected, as the statistical uncertainty scales as $1/\sqrt{N}$, where $N$ is the number of measurement shots.
+When the statistical error approaches or exceeds the magnitude of the systematic errors, the calibration cannot perform effectively.
+To model decoherence, we introduced bit-flip and phase-flip noise channels parameterized by mixing factors $p_{x}$ and $p_{z}$, respectively.
+These parameters range from 0 to 1, where 0 corresponds to no noise and 1 corresponds to complete randomization by the respective channel.
+We found that the method remains effective at moderate noise levels ($p_{x}=p_{z}=0.01$), and only degrades significantly at stronger noise ($p_{x}=p_{z}=0.1$), where its performance becomes comparable to the uncalibrated approach.
+See Supplementary Material S3 for details.
+In the multi-qubit setting, quantum state tomography is typically performed using local measurements on each qubit.
+In this case, our calibration method can be independently applied to each local measurement device.
+This approach enables the calibration to scale linearly in both measurement resources and computational cost, as each qubit requires only a fixed number of probe states and measurement settings (e.g., 30 states measured along six axes).
+Moreover, when qubits are independently addressable and correlated noise is negligible, calibration across qubits can be executed in parallel, effectively reducing the total measurement time to that required for a single qubit.
+The post-processing step, based on our optimization protocol, also parallelizes naturally across qubits.
+A more detailed discussion is provided in Supplementary Material S4, where we demonstrate through numerical simulations that calibrating individual qubits leads to improved reconstructions of $n$-qubit states.
+We also include a use example in the two-qubit photonic experiment [29].
+IV Special case: multiplicative error
 An important subclass of the problem is the case of multiplicative measurement operator error, where
-$\theta_{j}^{\prime}=(1+\delta)\theta_{j},$ (6)
-$\phi_{j}^{\prime}=(1+\epsilon)\phi_{j}.$ (7)
+$\tilde{\theta_{j}}=(1+\delta)\theta_{j},$ (6)
+$\tilde{\phi_{j}}=(1+\epsilon)\phi_{j}.$ (7)
 This case is relevant when we sequentially switch the unitary operations that precede a fixed projector to perform the tomographic measurements.
@@
 Because parameters $(\delta,\epsilon)$ are shared among all projection operators, the optimum search simplifies from 12 parameters to just two.
-In Section S2 of the Supplementary material, we show that with just eight probe states, the optimization landscape possesses a single minimum, and therefore standard gradient-descent optimization is applicable.
-The advantage is better performance when compared to global optimization and also the smaller probe state set.
-Another advantage is that the global optimum is now unique, and the method unambiguously finds the true set $\{\hat{\pi}_{j}\}$ even when there is an unknown unitary operation between preparation and tomographic measurements.
+In Supplementary material S5, we show that with just eight probe states, the optimization landscape possesses a single minimum, and therefore standard gradient-descent optimization is applicable.
+We observe that with the less general error model, one can tailor the probe state distribution to achieve the calibration with a smaller probe state ensemble using fewer computational resources.
+We further illustrate this on an edge example of the phase over-rotation model in Supplementary Material S5.
+Another advantage in the case of multiplicative errors is that the global optimum is now unique, and the method unambiguously finds the true set $\{\hat{\pi}_{j}\}$ even when there is an unknown unitary operation between preparation and tomographic measurements.
+And finally, one can take $\epsilon$ and $\delta$ into account when controlling $\theta_{j}$ and $\phi_{j}$ to achieve a more precise control of the projections.
+Having mitigated systematic errors in the measurement devices, we next extend the method to calibrated state preparation, solving the proverbial egg-chicken problem of characterizing quantum measurement without precisely known quantum states.
 Figure 2: Experimental demonstration for a polarization-encoded photonic qubit.
@@
 In both panels, the reconstructed states are represented on the Bloch sphere using Hammer projection, and their purity is color-coded.
+V State preparation
+The state preparation typically suffers from imperfections too.
+Here, we use the symmetry in the Born rule to formally exchange state preparation and measurement to utilize the previously introduced method to correct state preparation.
+Usually, the qubit is initialized in state $|0\rangle$, then turned into state $|\psi_{j}\rangle=\hat{U}_{j}|0\rangle$ by unitary evolution $\hat{U}_{j}$.
+For perfect preparation, $\hat{U}_{j}$ has to be perfectly known.
+In practice, there might be a slight discrepancy between the desired and the actual evolution.
+Our calibration reveals this discrepancy and brings information about the actual unitary.
+The tomographic projection is also typically realized by some unitary evolution $\hat{V}_{k}$ and subsequent projection, i.e., $|\pi_{k}\rangle=\hat{V}_{k}^{\dagger}|0\rangle$.
+The state tomogram of the probe state $|\psi_{j}\rangle$ consists of measured probabilities $|\langle\hat{\pi}_{k}|\psi_{j}\rangle|^{2}=|\langle 0|\hat{V}_{k}\hat{U}_{j}|0\rangle|^{2}$.
+Here, we assume that the operator $\hat{V}_{k}$ is precisely known.
+If not, the calibration procedure described above might be used to improve the knowledge.
+Let us now leverage the Born rule symmetry, i.e.
+$|\langle\hat{\pi}_{k}|\psi_{j}\rangle|^{2}=|\langle\hat{\psi}_{j}|\pi_{k}\rangle|^{2}$, to reverse our calibration from measurements to preparations.
+Physical reversal of the process consists of preparing states $\hat{V}_{k}|0\rangle$ and projecting them onto states $\hat{U}_{j}^{\dagger}|0\rangle$, obtaining tomogram elements $p_{kj}\propto|\langle 0|\hat{U}_{j}\hat{V}_{k}|0\rangle|^{2}$.
+We can analyze the tomograms to improve our knowledge of the measurement operator, i.e., $\hat{U}_{j}^{\dagger}|0\rangle\langle 0|\hat{U}_{j}$, related to state preparation in the non-reversed situation.
+The reversal can be realized in the experiment without its physical rearrangement, by performing projections onto the original probe states in the tomographic device and preparing states corresponding to the original tomographical projectors instead of the original probes.
+Either way, we gain information about $\hat{U}_{j}$, increasing the accuracy of state preparation.
+As an example, first recall the situation from Section IV.
+To calibrate six measurement operators, we applied them to eight probe states.
+These measurements were implemented by unitary rotations $V_{k}$, followed by projection onto the state $|0\rangle$.
+The calibration effectively revealed information about the $V_{k}$, i.e.
+parameters $\delta$ and $\epsilon$ which can be used to more precisely control the projections.
+Next, one would prepare six new probe states corresponding to the original projection $V_{k}|0\rangle$.
+These states carry information about systematic errors in the preparation.
+Each of these would be projected onto the original eight probe states $U_{j}|0\rangle$.
+The resulting tomograms contain readings corresponding to imprecise preparations measured with precise measurements.
+It could then be analyzed to gain information about the parameters of the preparation device.
+This information can be used to correct the state preparation, as we will demonstrate experimentally in the following section.
 VI Experiment: polarization state tomography
-We experimentally demonstrate the method in the waveplate-based polarization state tomography scenario, an example of a situation with multiplicative errors.
+We experimentally demonstrate the method in the waveplate-based polarization state tomography scenario.
 At the end of the process, we will improve our knowledge of the waveplate retardance and consequently mitigate the reconstruction artifact.
-The experimental setup is depicted in Fig.
-1(a).
+The experimental setup is depicted in Figure 2(a).
 We used a continuous-wave 810 nm thermally-stabilized fiber-coupled laser diode (LD).
@@
 The deviation from the ideal retardance will introduce errors in the preparation.
-We assumed the perfect retardance in the first step of the experiment as our method does not require exact knowledge of probe states.
-To demonstrate the tolerance of the method to a unitary operation between preparation and tomographic measurement, we add a quarter-wave plate (QWPX) with its fast axis oriented at -12.5 deg relative to the horizontal direction.
+We assumed the perfect retardance in the state preparation stage as our method does not require exact knowledge of probe states.
+To demonstrate the tolerance of the method to a unitary operation between the preparation and tomographic measurement, we add a quarter-wave plate (QWPX) with its fast axis oriented at -12.5 deg relative to the horizontal direction.
 Because this unitary operation can be in principle unknown, we restrict ourselves from using any information about QWPX in the calibration.
@@
 The waveplate settings for the Pauli tomography are listed in Tab.
-S2 of the Supplementary Material.
-All waveplates were mounted into a precision motorized rotation stage that provided bidirectional repeatability of ±$\pm 30$ mdeg.
+S3 of the Supplementary Material.
+All waveplates were mounted into a precision motorized rotation stage that provided bidirectional repeatability of $\pm 30$ mdeg.
 We used reverse-biased PIN photodiodes (PD) to measure optical intensity at both outputs of the beam-displacer.
-We normalized the readings using the sum of two signals to eliminate laser power fluctuation.
+We normalized the readings using the sum of the two signals to eliminate the influence of laser power fluctuation.
 The action of the waveplate could be seen as the rotation of the Bloch sphere around axis $(\sin(2\alpha),0,\cos(2\alpha))$ with rotation angle $\Gamma$, where $\alpha$ determines the angular position of the waveplate and $\Gamma$ its retardance.
-The retardance usually slightly deviates from its nominal value.
+The retardance usually deviates from its nominal value.
 These deviations manifest as under- or over-rotation of the Bloch sphere.
-We parameterize the measurement operator by precisely controlled waveplate angular positions $x$, $y$, and half- and quarter-wave plate retardance deviations $\delta$ and $\epsilon$, respectively.
-These deviations are initially unknown to us, and we will use the presented method to estimate their values.
-We use eight probe states and perform the Pauli tomography for each.
-Initially, we assume zero deviations, $\delta=\epsilon=0$, and observe the purity modulation $\Delta P=0.0572(3)$.
-The number in parenthesis is one standard deviation at the last significant digit, e.g.
-$0.0572(3)=(572\pm 3)\cdot 10^{-4}$.
+So, in the context of the previous sections, the deviation of the effective wave plate retardance from its nominal value is the origin of the mismatch between the measurement operators.
+The retardance mismatch can be attributed to the manufacturing tolerances or to the use of a light at a slightly different wavelength than the nominal design value of the wave plate.
+The exact values of these deviations are initially unknown to us, and we will use the presented method to estimate their values more precisely.
+We parameterize the measurement operator by precisely controlled waveplate angular positions $x$, $y$, and half- and quarter-wave plate retardance deviations $\delta^{\prime}$ and $\epsilon^{\prime}$, respectively.
+We sequentially prepare eight distinct polarization states and perform the Pauli tomography for each.
+As we introduced in previous sections, these probe states have similar purity and are quasi-uniformly distributed over the Bloch sphere.
+We use them to probe the measurement device.
+Initially, we assume zero deviations, $\delta^{\prime}=\epsilon^{\prime}=0$, and observe the purity modulation $\Delta P=0.0572(3)$.
+The number in parentheses is one standard deviation at the last significant digit, e.g.
+$0.0572(3)=(572\pm 3)\times 10^{-4}$.
 The uncertainties were estimated using bootstrapping with the assumption of normally distributed photo-current with 0.1% relative standard deviation, which is the most dominant term and is related to the ammeter precision.
 The shot noise is negligible at the power level of our laser, compared to the technical noise.
-Then, we vary parameters $\delta$ and $\epsilon$ and update the measurement operators employed in the reconstruction to minimize the purity modulation.
+Then, we vary parameters $\delta^{\prime}$ and $\epsilon^{\prime}$ and update the measurement operators employed in the reconstruction to minimize the purity modulation.
 The optimal $\delta$ and $\epsilon$ are close to the true values because the reconstruction artifact is significantly reduced to $\Delta P=0.0140(2)$.
-We found the optimal parameters $\delta=4.5(2)$ deg, $\epsilon=-1.30(4)$ deg.
-We then reverse the process to determine retardance deviations $\tilde{\delta},\tilde{\epsilon}$ of the waveplates used in state preparation.
+We found the optimal parameters to be $\delta^{\prime}=4.5(2)$ deg, $\epsilon^{\prime}=-1.30(4)$ deg.
+We then reverse the process to determine retardance deviations $\tilde{\delta},\tilde{\epsilon}$ of the waveplates used in the preparation stage.
 We prepare the six eigenstates of Pauli operators, project each onto eight states corresponding to the original probe state set, and treat this data as before.
@@
 To reconstruct the quantum states, we could directly use this knowledge to update our measurement operators.
-However, we take one extra step which allows us to prepare the states with greater precision and control the measurement operators precisely.
+However, we take one extra step, which allows us to prepare the states with greater precision and control the measurement operators precisely.
 We took these deviations into account in the verification step and calculated new angular positions of wave plates for tomographic projections and state preparation.
@@
 As a reference, we first assumed perfect waveplates and performed state tomography for each test state.
-We indeed observed purity modulation $\Delta P=0.0607(2)$ with purity minimum $P_{\min}=0.9383(2)$.
+We observed purity modulation $\Delta P=0.0607(2)$ with purity minimum $P_{\min}=0.9383(2)$.
 This modulation indicates reconstruction artifacts.
@@
 The information gained in the calibration clearly improved the reconstruction.
+The overall fidelity $F=0.997(3)$ surpasses the value 0.99(1) reported previously by [15].
+Note that this fidelity is calculated relatively to the expected ideal state and was evaluated on a set of 58 test states, independently of calibration.
+This is in contrast to the previous demonstration, where only 14 states were used, and these states were also used for calibration.
 The results in Figure 2 indicate improved quality of state preparation and tomographic measurements when we use the knowledge of $\delta$ and $\epsilon$ to improve both preparation and measurement.
 The small remaining artifacts are present due to the limited precision of the learned retardance deviations and also the repeatability of the wave plate rotation.
+Having demonstrated experimental feasibility with polarization-encoded optical qubits, we now turn to on-chip photonic qubits and illustrate how our calibration method applies to path-encoded qubits in this increasingly relevant setting.
 VII Path-encoded qubits on a photonic chip
-Figure 3: Application of the method for on-chip photonic path qubit.
+Figure 3: Application of the method for an on-chip photonic path qubit.
 (a) Projection apparatus for a path qubit.
@@
 The photons in the upper output path are detected with a single-photon detector (SPAD).
-(b) 1D-section of 6D function $\Delta P(c_{1},\dots,c_{6})$.
-The value of $\Delta P$ at the initial assumption of $c_{j}$ is $1.8\cdot 10^{-2}$ on the ensemble of probe states.
-Minimum $\Delta P=4.5\cdot 10^{-4}$ lies at the true value $c^{\prime}_{6}$.
-(c) Initial mismatch results in state-dependent purity artifact.
+(b) 1D-section of 6D function $\Delta P(c^{\prime}_{1},\dots,c^{\prime}_{6})$.
+The value of $\Delta P$ at the initial assumption $c^{\prime}_{j}$ is $1.8\times 10^{-2}$ on the ensemble of probe states.
+Minimum $\Delta P=4.5\times 10^{-4}$ lies at the true value $c_{6}$.
+(c) Initial mismatch results in a state-dependent purity artifact.
 We plot test states on the Bloch sphere as circles and color-code their reconstructed purity.
 (d) When we consider the true retardance, the artifact vanishes.
-As another example, we numerically illustrate the measurement-device agnostic calibration for tomography of photonic on-chip path qubits [30, 35, 36, 37].
-The qubit is encoded into the photon propagation in one or another waveguide (states $\ket{\uparrow}$ or $\ket{\downarrow}$), as depicted in Figure 3(a).
-The local heating of one waveguide controls the phase between the two paths.
-The heating is typically achieved by an electrical current flowing through a small heater element.
-A Mach-Zehnder interferometer serves as a variable coupler, and its interferometric phase controls the coupling ratio.
-Detection of photons in a single optical path implements a projective measurement of the path qubit.
-The circuit in Figure 3(a) implements a projection onto state |$\ket{\pi}$, whose colatitude and longitude on the Bloch sphere are controlled by optical phases.
-The relation between optical phases $\varphi_{1,2}$ and voltages $V_{1,2}$ applied on the heating elements is approximated with polynomials [35]
+[…]
 $\varphi_{1}(V_{1})=c_{1}+c_{2}V_{1}+c_{3}V_{1}^{2},$ (13)
@@
 3(c), shows the measurement operator mismatch artifact.
-We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j})$ where the systematic error is mitigated.
-Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c^{\prime}_{6}$.
-Taking the true values $c^{\prime}_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).
+We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j}^{\prime})$ where the systematic error is mitigated.
+Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c_{6}$.
+Taking the true values $c_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).
 VIII Conclusion
@@
 The application of the reported measurement-device agnostic quantum measurement goes beyond the full quantum tomography [11, 22].
-The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches such as Monte-Carlo sampling [38, 39, 40, 41], permutationally invariant tomography [42], compressed sensing [43, 44], etc.
+The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches, such as Monte-Carlo sampling [39, 40, 41, 42], permutationally invariant tomography [43], compressed sensing [44, 45], etc.
 The method can also improve the accuracy of building and characterization of the photonic and quantum circuits in complex media [46, 47].
@@
 B.
-acknowledges the support of Palacký University under grant No.
-IGA-PrF-2024-008.
+acknowledges the support of Palacký University under grants No.
+IGA-PrF-2024-008 and IGA-PrF-2025-010.
 We acknowledge the use of cluster computing resources provided by the Department of Optics, Palacký University Olomouc.